### v3

# How Good is ChatGPT in Giving Advice on Your Visualization Design?

Abstract.
Data visualization creators often lack formal training, resulting in a knowledge gap in design practice. Large-language models like ChatGPT, with their vast internet-scale training data, offer transformative potential in addressing this gap. In this work, we first utilized the VisGuide forum, a repository of data visualization questions, to compare ChatGPT-generated responses with Human replies. We additionally conducted a user study delving into practitioners’ reactions and attitudes toward ChatGPT as a visualization assistant. Participants, who brought their visualizations and questions, received feedback from both Human experts and ChatGPT in a randomized order. The results highlight the unique advantages and disadvantages of ChatGPT, such as its ability to quickly provide a wide range of design options based on a broad knowledge base, while also revealing its limitations in terms of contextual understanding, interaction dynamics, and perceived trustworthiness of responses. From these insights, we derive design considerations for LLM-based design feedback systems.
1. Introduction
Visualizations are ubiquitous and widely employed by practitioners across various disciplines. However, many of those creating visualizations commonly lack formal training in visualization design and instead acquire the necessary skills on the go, leaning from examples from the internet, online blogs, and other publicly available resources (Esteves and Neves, 2022; dvs, [n. d.]). These people may find themselves in situations where they must navigate intricate design choices (Choi et al., 2023b) and consequently revert to their instincts, drawing from the experiences and observations they have accumulated along the way (Choi et al., 2023b; Bako et al., 2023). Alternatively, they seek feedback from colleagues who can help provide novel perspectives, validate design choices, and challenge assumptions (Choi et al., 2023b; Luther et al., 2015). Nevertheless, not all practitioners have the privilege of accessing such feedback from experienced colleagues.
This paper delves into ways to address this knowledge gap in practical settings. We examine the capabilities of large language models (LLMs) equipped with extensive internet-scale training data, which holds transformative potential (Okerlund et al., 2022). LLM-based chatbots, exemplified by ChatGPT (OpenAI, [n. d.]), can offer tailored suggestions by analyzing a practitioner’s design choices and considering established best practices. This emerging advancement can assist casual and novice practitioners, who may have limited knowledge and resources in data visualization, by providing human-like guidance throughout their design journey.
However, open questions remain regarding the extent of LLMs’ current knowledge of visualization best practices, the quality and practical applicability of their suggestions, and how these suggestions differ from those provided by human expert visualization designers. Addressing these open questions, we have formulated two specific research questions to guide our investigation:
RQ1. Can ChatGPT rival Human expertise in data visualization knowledge?
RQ2. How would visualization practitioners perceive ChatGPT’s design feedback?
We employ a mixed-methods approach (Fig. 1) to explore the potential of ChatGPT in offering data visualization design knowledge. We selected ChatGPT as a representative proxy for LLMs due to its widespread use and generally superior performance compared to other LLMs (Borji and Mohammadian, 2023; Ahmed et al., 2023; Giannakopoulos et al., 2023; Lim et al., 2023). First, we investigated posts and discussions at the VisGuides forum (Diehl et al., 2018), which captures public questions and responses about visualization design. We asked the questions to ChatGPT and compared its responses to the Human responses on VisGuides across six key metrics: coverage, topicality, breadth, clarity, depth, and actionability. We found that ChatGPT’s performance is comparable to, and often better than, Human responses, especially in providing clear and comprehensive answers that cover all questions asked. In contrast, human responses typically offered specific suggestions, like citing references beyond basic improvements, but showed greater variability due to multiple respondents involved.
While the VisGuide analysis provided insight into ChatGPT’s response quality (RQ1), it fell short of capturing practitioners’ reactions, attitudes, and overall experiences (RQ2). Thus, we conducted a qualitative user study wherein we invited practitioners to partake in a comparative assessment of Human expert feedback and ChatGPT-generated feedback. For this study, participants were invited to bring their own visualization designs and projects to our sessions for questions and feedback. Interestingly, all participants preferred discussing with Human experts, valuing their ability to engage in fluid conversations and provide tailored recommendations. However, participants appreciated the broad knowledge of ChatGPT, particularly its ability for quick brainstorming and ideation from a neutral standpoint, while acknowledging its limitations in text-based communication.
Both studies consistently shed light on the strengths and weaknesses of ChatGPT and Humans. ChatGPT excelled at providing broad suggestions, whereas Humans demonstrated a better grasp of nuanced details for more in-depth ideas. A stronger preference for Humans in the feedback study was linked to the more consistent opinions and natural conversational dynamics of Human experts. These findings also point to areas for future enhancement of LLMs, including enhancing visual perception and proactive discussion capabilities to better tackle the nuanced contexts of design problems. Furthermore, we discuss the potential role of an LLM-based knowledge companion in visualization practice and, more broadly, in education.
Our work represents a meaningful step forward for the field amidst the rise of LLM-driven visualization innovations. The contributions of this work lies in the following:
An evaluation framework for assessing the knowledge of LLMs in data visualization design, based on metrics derived from the analysis of the VisGuides forum.
Insights into LLM-based chatbots in visualization practice via the lens of ChatGPT analysis to highlight their strengths and weaknesses compared to Human counterparts.
Future considerations for enhancing the user interface of LLM-based chatbots as design companions, aiming for improved human-AI interaction.
2. Related Work
2.1. Data Visualization Design Practice
Data visualization has gone mainstream, frequently used to explore and communicate data in a wide variety of industry sectors. As a result, data visualization practitioners have diverse professions, such as managers, scientists, journalists, and analysts, not necessarily having a specific job title as “visualization designer” or “visualization developer.” (Esteves and Neves, 2022) These professionals do not typically have formal training or education in data visualization, while mostly learning necessary visualization skills on the go (Esteves and Neves, 2022; dvs, [n. d.]). While decades of empirical research in data visualization have produced fruitful knowledge in effective visualization design, it is unclear how well it is accessible to practice (Choi et al., 2021; Diehl et al., 2018). In fact, it has been reported that many practitioners are not aware of theories and principles from research (Parsons, 2022).
As with any creative design profession, visualization practitioners run into challenging situations where they must make decisions among conflicting design alternatives (Choi et al., 2023b). Such decisions might include selecting chart types, making competing decisions between aesthetics and functions, picking appropriate color scales, and addressing conflicts with user and business stakeholder needs (Choi et al., 2023b). Processes by practitioners to address such design issues are more nuanced and multi-faceted than theoretical process models proposed in the research field (Parsons, 2022). They depend on a kind of situated planning that often relies on forms of experimentation and responding to what is happening in the moment (Parsons, 2022; Parsons et al., 2020). They balance readability and engagement and also consider contextual factors such as whether their visualizations are business-oriented dashboards or public-facing graphics (Parsons and Shukla, 2020).
Practitioners frequently resort to their intuition or gut instinct while also looking for examples for inspiration (Choi et al., 2023b; Parsons, 2022). Moreover, they also seek feedback from their colleagues to improve their data visualizations (Choi et al., 2023b; Esteves and Neves, 2022). Such feedback is essential to assess a design and generate revision ideas in the design process, although the fear of criticism and non-anonymity can make people uncomfortable receiving feedback from colleagues. Moreover, self-employed people might struggle to find such colleagues who can provide valuable feedback (Chalkidis et al., 2019)
Previous research has utilized online crowdsourcing to streamline the feedback collection process, providing rapid access to large, diverse participant pools and enabling faster feedback cycles at reduced costs (Buhrmester et al., 2011; Haug and Mädche, 2021; Yen et al., 2016; Luther et al., 2015). To address the challenge of insufficient expertise among crowd workers, researchers have explored various techniques to enhance feedback quality, including micro-tasking (Xu et al., 2014; Luther et al., 2014), implementing predefined rubrics (Yuan et al., 2016), providing illustrative examples (Kang et al., 2018), and establishing feedback guidelines (Krause et al., 2017). While these efforts have primarily focused on feedback for graphic design projects, recent studies in data visualization have investigated automated methods for providing perceptual feedback through image processing techniques (Shin et al., 2023) and developing structured methods for collecting empirically-driven quantitative feedback (Choi et al., 2023b).
In this work, we examine how an LLM-based chatbot, trained with extensive knowledge in the wild, can serve as a design companion to provide knowledge and feedback to varying backgrounds of practitioners. Can LLMs replace the feedback gathering process? We aim to bridge the gap between LLMs’ theoretical potential and their practical application in data visualization, a critical endeavor in the evolving landscape of artificial intelligence within the field.
2.2. Evaluating LLMs’ Knowledge Capacity
LLMs, such as BERT (Devlin et al., 2018), GPT (Brown et al., 2020; OpenAI, 2023), and Llama (Touvron et al., 2023), have rapidly gained popularity and adoption across both industry and academia (Bommasani et al., 2021; Zhao et al., 2023). Leveraging massive amounts of text data and computational power, these foundation models have achieved impressive performance on a wide variety of tasks, from generating human-like text to high-quality images (Radford et al., 2019, 2018; Wei et al., 2022a). However, the scale and complexity of these models make it difficult to fully understand the scope and limitations of their language capabilities (Bender et al., 2021).
A number of benchmark tasks have been proposed to assess the general knowledge and reasoning capabilities of language models in a more holistic manner (Liang et al., 2022; Chang et al., 2023). These include tasks that require models to answer broad-coverage factoid questions based on pre-existing knowledge (Liang et al., 2022). Other benchmarks evaluate numeric, analogical, and logical reasoning skills across different modalities (Qin et al., 2023; Zhang et al., 2023b), while some others examine the ability of LLMs to generalize knowledge to new concepts (Wei et al., 2022a). Assessing models on such diverse reasoning skills can provide deeper insight into their adaptability, knowledge gaps, and limitations. In addition to assessing such capabilities, an important area of research has focused on evaluating potential harms and biases in large language models. Some studies have performed targeted adversarial evaluations to expose problematic biases encoded in the models’ training data (Nadeem et al., 2021; Tan and Celis, 2019; Gehman et al., 2020).
In addition to generic capabilities, research has been carried out to investigate LLMs’ discipline-specific knowledge and skills. The inherent intricacies and nuances of specific domains necessitate tailored evaluation methods. For instance, researchers explored students’ and teachers’ usage of and perceptions toward LLMs, investigating opportunities and potential benefits for education (Kasneci et al., 2023; Baidoo-Anu and Owusu Ansah, 2023; Tlili et al., 2023). Researchers also tested LLMs’ ability to pass exams in specific domains like medicine (Gilson et al., 2023), law (Fei et al., 2023; Choi et al., 2023a), and business (Mollman, 2023). Others have evaluated LLMs’ ability to carry out real-world tasks. Examples include analyzing diagnostic abilities by answering questions about medications, symptoms and treatments (Lee et al., 2023); producing research hypotheses (Lahat et al., 2023); predicting legal judgment and recommending case citations (Chalkidis et al., 2019); and predicting chemical properties (Pan, 2023).
Recent studies have initiated the exploration of Large Language Models’ (LLMs) applicability in the realm of data visualization. Some investigations have examined their capacity to generate data visualizations (Paik, 2023; Hutchinson et al., 2023), facilitate exploratory data analysis (Dibia, 2023), and create stylistic data graphics (Dibia, 2023), in addition to generating graph drawings (Bartolomeo et al., 2023). Furthermore, studies have delved into the usability of LLM-based code generation in coding data visualizations (Vaithilingam et al., 2022), as well as understanding visualization practitioners’ perceptions regarding generative text-to-image models (Schetinger et al., 2023).
Unlike the previous work, we focus on evaluating the design knowledge capacity of LLMs. This facet is one of the primary pillars of data visualization skills, extending beyond the mere creation of data visualizations. In other words, this work delves into the cognitive aspects of LLMs, assessing not just their visualization code output but the reasoning and principles underlying their suggestions.
3. Assessing ChatGPT’s Competence in Data Visualization Knowledge
As a step toward assessing ChatGPT’s data visualization knowledge, we were interested in its ability to respond to real-world data visualization questions in comparison to Humans.
3.1. Data Collection
We used VisGuides (Diehl et al., 2018) as our main source for visualization questions. VisGuides is a discussion forum focused on visualization guidelines, where practitioners from various backgrounds, such as scientists, designers, engineers, and students, ask and respond to questions regarding their visualizations and fundamental visualization principles.
We identified a total of 226 questions within the VisGuides repository. Among these, we selected 119 questions for the study based on specific inclusion criteria. These criteria included the presence of a question in the post, the requirement for a visualization description to possess sufficient visual encoding (due to ChatGPT’s inability to process images at the time of our analysis), the need for questions to be comprehensible and well-structured (excluding overly generic inquiries like ’What do you think?’ or ’Is this good enough?’), and the necessity of having at least one Human response. Two of the authors jointly reviewed and unanimously approved the inclusion of these 119 questions.
We discovered that Visguides primarily featured two types of questions, broadly divided into two categories: design feedback questions (87) and visualization guideline questions (32). The former aimed to enhance users’ personal visualizations (see Figure 3 B & C), while the latter sought to comprehend visualization guidelines or principles (see Figure 3 A). For a more comprehensive taxonomy of questions on the platform, Diehl et al. provide detailed insights (Diehl et al., 2021).
After compiling the list of qualifying questions, we presented each query to ChatGPT in May 2023, accompanied by a role-playing prompt designed as follows:
Please act as if you are a visualization expert.
Can you respond to this question delimited by ///.
{Question}
To streamline our analysis, we created a spreadsheet where each row included a question, the corresponding ChatGPT response (GPT 3.5), and a compilation of Human responses. Each question had 1.54 Human responses on average ( = 1.03). The average approximate word count of the Human responses was 235.35 ( = 289.62) with a maximum of 1721 words and minimum of 13 words. The average approximate word count of the ChatGPT responses was 463.42 ( = 189.34) with a maximum of 1192 words and minimum of 115 words.
3.2. Analysis Decisions
Why Did We Use VisGuides?
We explored other sources of data visualization questions in the wild, including StackOverflow, the #help-general channel in the Slack workspace of The Data Visualization Society (dat, [n. d.]), and r/dataisbeautiful on Reddit. StackOverflow mainly focuses coding-related questions, which differ from our focus on design knowledge questions. The Data Visualization Society’s Slack channel, limited by a lack of a paid license, does not archive questions older than 90 days, and the queries are mostly technical, such as using specific software tools to create charts or visual effects. Reddit’s r/dataisbeautiful channel aligns more closely with our objectives, but a thorough review revealed that it contains more discussions and occasional critiques of web-found charts than actual questions. Hence, we decided to focus on the questions in the VisGuides forum.
Why Did We Choose ChatGPT Over Other Chatbots?
We selected ChatGPT as a representative of LLM-based chatbots for several reasons. While these chatbots use the same transformer architectures, ChatGPT is the most widely adopted in practice. Also, in our early pilot experimentation with Bard (now Gemini), Claude, and Bing (which uses the same GPT model as ChatGPT), ChatGPT generally provided responses that were better aligned in terms of style and content. For example, Bing often veered off by drawing information from various web sources, which were frequently inferior to ChatGPT’s internal knowledge responses. Furthermore, Bard’s recommendations were narrower in scope and often not applicable, like suggesting the addition of a trend line to a Gantt Chart. Overall, our impression was that these systems are mostly comparable, with ChatGPT having a slight edge over the others. Several studies also indicated that ChatGPT demonstrates superior performance compared to other LLMs in terms of a variety of task questions involving reasoning, facts, and fairness (Borji and Mohammadian, 2023; Ahmed et al., 2023). That said, we plan to establish a comparative benchmark which we discuss in the future work section.
Why Did We Not Use Different Prompting Methods?
ChatGPT and other LLM-based chatbots are designed to follow human instructions (Wu et al., 2023; Liu et al., 2021b). Different prompting techniques, such as chain-of-thought (Wei et al., 2022b) and self-consistency (Wang et al., 2022), can lead to varied responses. To ensure a more objective evaluation, we used neutral prompts without any specific instructions (e.g., making the response concise) other than the role prompting. Moreover, we assume that users typically would not employ complex prompting strategies and would most likely use zero-shot queries. Thus, we merely forwarded the question from VisGuides to ChatGPT.
3.3. Analysis Method
Two researchers performed an open coding process on a randomly selected 20% sample of the questions and the corresponding ChatGPT and Human responses. This process resulted in the development of six evaluation metrics: breadth, clarity, depth, actionability, coverage, and topicality.
These metrics were assessed using a Likert scale that ranged from 1 to 5. For instance, a rating of 1 indicated narrow responses, while a rating of 5 denoted very broad responses in the context of the breadth metric. The researchers initially assigned scores to the 20% sample, engaging in discussions to resolve any discrepancies and establish consensus definitions and criteria for each score within every metric to ensure consistency. Following this, one of the researchers proceeded to score the remaining question responses. Throughout this coding process, they held continuous meetings to discuss any new issues and questions to ensure consistency and quality of codes and scores, refining metric definitions and score applicability.
Given that some questions elicited multiple Human responses while others received only one, We created a composite score to evaluate the collective quality of the Human responses. Furthermore, any elements of the Human responses that extended beyond the capabilities of ChatGPT v3.5, such as providing citations to related literature, links to external resources, visual examples, or posing follow-up questions, were recorded. We also categorized whether each question sought feedback on the questioner’s visualizations or inquired about general visualization design knowledge.
3.4. Results
3.4.1. Ratings on Quality Metrics
Figure 2 shows the scores of response quality metrics for both ChatGPT and Human responses. We describe the findings for each quality metric as below.
Coverage
Both ChatGPT and Humans excelled in covering the entirety of the question, with ChatGPT slightly outperforming Humans. ChatGPT scored 4.92 ( = 0.33), while Humans scored 4.37 ( = 0.95). The lower score often stemmed from the Human respondent’s failure to address all facets of the question. For instance, in Figure 3C, the user raised a two-part question, “…would you recommend all the states be placed on the Y-axis instead of grouped as regions? If so, which visual layouts would be effective with dealing with the large number of states?…”. However, the Human response only addressed the second part of the question, suggesting a heat map as a useful visual layout.
Topicality
This score indicates the degree to which the response stays on topic. ChatGPT averaged 4.87 ( = 0.45), while Humans averaged 4.72 ( = 0.59). There were moments when both ChatGPT and Humans strayed from the main topic. For instance, when a user inquired about the application of visual variables in VR visualizations,111visguides:visual-variables-in-vr ChatGPT not only addressed this question but also came up with a hypothetical question asking what factors to consider when choosing such visual variables. Human respondents also occasionally veered off-topic by offering unsolicited advice and extending their responses beyond the original question’s scope.
Breadth
Breadth measures how widely a response explores various ideas, concepts, options, or perspectives. In this category, ChatGPT outperformed Human responses by a significant margin. ChatGPT achieved an average score of 4.62 ( = 0.61), while Human responses averaged only 3.35 ( = 1.13). For example, in Figure 3A, when the user inquired about color palette guidelines, ChatGPT provided an extensive list of potential approaches, encompassing data type, color harmony, accessibility, etc. In contrast, the Human respondent to this question concentrated primarily on the creation of domain-specific color maps, offering in-depth insights into the process.
Clarity
The clarity score assesses how easily a reader can comprehend a response, taking into account conciseness, lack of verbosity, and the organization of the response structure. Notably, ChatGPT consistently presented its responses in a well-structured list. This contributed to ChatGPT receiving a high clarity score, averaging 4.93 ( = 0.28). Human responses also demonstrated good clarity with an average score of 4.30 ( = 0.79). They received lower ratings when the responses were overly brief, making it difficult to discern the specific reference points.
Depth
The depth score evaluates the extent of explanations, expertise, and valuable insights in a response. Human responses averaged a score of 3.52 ( = 1.07), while ChatGPT’s average was 3.44 ( = 0.71). Although the scores are comparable to each other, a distinct difference in depth between a ChatGPT response and a Human response is evident in Figure 3B. The ChatGPT response primarily focuses on general, surface-level improvements applicable to any visualization, such as adding labels and legends. In contrast, the Human response often provides more tailored and specific suggestions for enhancing the visualization, as demonstrated by the recommendation to “put a label before the numerical attribute like ’Estimated Production: 1000’.” This detailed guidance directly pertains to the unique aspects of the visualization in question. However, Human responses also displayed greater variability. For instance, when asked about color clarity and alternative data representation to reduce clutter, a Human response simply confirmed the current design choices are fine222visguides:latitude-distribution-of-solar-and-wind-farms-uk. The lower scores of such poor quality single responses pulled the average down.
Actionability
The actionability score assesses whether the response offers guidance that can be readily implemented in the visualization. This metric is not applicable to visualization guideline questions. The actionability scores were quite similar: Human responses had an average score of 3.86 ( = 0.94), while ChatGPT achieved an average of 3.99 ( = 0.61). Overall, ChatGPT received a high score for actionability, as it consistently provided a list of recommendations. However, the score was occasionally reduced due to the generic or vague nature of these suggestions. For instance, in Figure 3B, ChatGPT offered advice for addressing label clutter, selecting colors, and refining the legend. Nevertheless, these recommendations were explained in a manner that made them applicable to most visualizations. The actionability scores for Human responses tend to be low when the Human refrains from offering recommendations and instead validates all of the user’s choices333visguides:comparison-of-acute-hepatitis-from-tycho-dataset.
3.4.2. Further Observations in Human and ChatGPT Responses
Humans tend to provide external resources
—43 out of 119 (36.13%). Human responses included elements that ChatGPT would not provide unless explicitly instructed; even when asked, it is known to frequently provide incorrect citations (Frosolini et al., 2023). This additional content encompassed references to academic research, links to related articles and websites, citations of studies, and the inclusion of informative video links. For instance, in Figure 3A, asking for guidelines on choosing color palette, the responses include references to three different academic research papers including “H. Fang, et al.”. In another case, when the question was about using blow-apart effects, the Human respondent embedded an educational video related to the subject444visguides:the-blow-apart-effect.
Multiple Human responses tend to complement each other
—Most questions had a single response (84), while 35 questions had multiple responses. Average ratings generally favored responses with multiple contributors across all categories, except for topicality. Significant rating differences were observed in breadth, depth, and actionability. Single-respondent questions averaged 3.06 ( = 1.05) in breadth, 3.35 ( = 1.05) in depth, and 3.72 ( = 0.96) in actionability. Conversely, multi-respondent questions scored higher, averaging 4.06 ( = 0.98) in breadth, 3.94 ( = 0.98) in depth, and 4.26 ( = 0.74) in actionability. When there’s only one response to a question, the quality of that response becomes heavily reliant on the individual respondent, leading to significant variability in quality, as seen in Figure 3A (high quality) versus Figure 3C (low quality). However, when multiple respondents contribute, they complement each other and compensate for areas in which one respondent might fall short (Figure 3B).
Human response rating improves for visualization guideline questions
—For ChatGPT responses, the differences in ratings between feedback and visualization guideline questions consistently remained small (less than 0.2 rating points). However, Human responses received higher scores for addressing guideline questions and lower scores for addressing design feedback questions across all metrics. For instance, Human responses excelled in breadth and depth for visualization guideline questions, with average ratings of 4.09 ( = 0.95) for breadth and 4.25 ( = 0.97) for depth. In contrast, design feedback questions received lower average ratings, with 3.08 ( = 1.06) for breadth and 3.25 ( = 0.97) for depth.
Guideline questions frequently require broader perspectives and readily accessible design knowledge, with 84.4% (27/32) of their responses citing external references, in contrast to design feedback questions, where only 18.39% (16/87) of Human responses did so. Conversely, feedback questions entail an understanding of domain-specific data and tasks, potentially making it challenging to offer comprehensive insights. For instance, a question like “Is this color scheme suitable?” requires a deep understanding of the domain. Conversely, questions such as the one about gendered colors in visualizations555visguides:use-of-gendered-colours-in-visualization can easily elicit a variety of various viewpoints and existing resources.
Question specificity contributes to response variability
—While we did not quantify this, we observed ChatGPT’s response might be susceptible to how specific and clear the question was. Figure 3B shows such an example when the user query was “do you believe the graph is clear?” Other examples of less specific user questions include: “is my color map optimal?” and “how can my visual design be improved?”666visguides:map-and-bar-visualization. Such commonly encountered ill-specified questions can pose a challenge for LLMs, as they may begin suggesting irrelevant feedback instead of asking clarifying questions. On the other hand, in Figure 3C, the question is more specific, resulting in more meaningful options and insights offered by ChatGPT.
ChatGPT’s color vision deficiency is usually not problematic—Color-related questions were prominent. Out of the 119 questions, 32 were centered on color-based design feedback (e.g., “Is the choice of the colour scheme appropriate?” 777visguides:area-chart-generation-capacity-over-the-years). Users rarely mentioned the specific colors used in their visualizations, often providing only visual encoding information. As a result, ChatGPT’s responses were typically based on general color principles. Despite this limitation, ChatGPT surpassed its overall average in most categories, except for depth and actionability.
3.5. Takeaways
Our analysis reveals ChatGPT’s ability to respond to data visualization queries compared to Human counterparts. ChatGPT excelled in breadth, exploring a wide range of ideas and concepts, surpassing Human responses. Clarity was another area where ChatGPT scored highly, benefiting from its structured response format. However, ChatGPT’s suggestions often leaned towards being more generic. In contrast, Human experts often offered more tailored and specific advice; they enhanced their responses by providing reference resources and complementing each other in multi-respondent scenarios. While these findings are promising, they do not provide insights into how practitioners perceive the value of Human’s responses, prompting the need for further research.
| Participant ID | Role | Years in Data Vis | Gender | Age range | Racial background | Frequency of ChatGPT usage |
| P10 | Developer | 3-5 years | Male | 25-34 years old | Asian | Daily |
| P11 | Manager | 1-3 years | Male | 25-34 years old | Caucasian | Occasionally |
| P12 | Freelancer | 3-5 years | Male | 35-44 years old | Asian | Daily |
| P13 | Journalist | 1-3 years | Female | 18-24 years old | Hispanic | Weekly |
| P14 | Consultant | 3-5 years | Male | 25-34 years old | Caucasian | Daily |
| P15 | Product Designer | 3-5 years | Female | 25-34 years old | Caucasian | Weekly |
| P16 | Analyst | 5-10 years | Male | 35-44 years old | Caucasian, African American | Weekly |
| P17 | Scientist | 3-5 years | Female | 25-34 years old | Asian | Weekly |
| P18 | Scientist | 1-3 years | Female | 25-34 years old | African American | Weekly |
| P19 | Student | 5-10 years | Male | 25-34 years old | Caucasian | Only once or twice |
| P20 | Freelancer | ¿ 10 years | Male | 45-54 years old | Caucasian | Occasionally |
| P21 | Student | 3-5 years | Male | 25-34 years old | Asian | Daily |
4. Perception of practitioners toward ChatGPT’s utility
Building upon our previous research, we conducted a comparative interview study involving feedback sessions with ChatGPT and Human experts. We aimed to gain a deeper understanding of how data visualization practitioners perceive ChatGPT’s role as a design companion, uncovering their attitudes and perceptions regarding ChatGPT’s design feedback.
4.1. Recruitment
We aimed to engage a diverse group of data visualization practitioners in our study. We recruited participants through academic mailing lists targeting students and scientists, as well as the Data Visualization Society’s Slack channel (dat, [n. d.]). Inclusion criteria required proficiency in English, experience in data visualization creation, and a willingness to share at least one of their data visualizations. Initially, 41 individuals responded to the recruitment survey, from which we selected 12 participants to take part in the study.
4.2. Participants
The study involved 12 participants (P10 to P21), creating a diverse group with varying professional backgrounds, experience levels, and familiarity with ChatGPT (see Table 1). Their roles encompassed a broad spectrum, including developers, managers, journalists, consultants, product designers, analysts, scientists, and students. Professional experience ranged from less than one year to over ten years in their respective fields. Furthermore, participants’ engagement with ChatGPT varied, with some being daily users while others interacted with it only occasionally or as the need arose.
4.3. Tasks & Procedures
Before each study session, participants were asked to prepare a visualization they were comfortable sharing, along with a list of relevant questions. These sessions were conducted using Zoom between the end of July and early August 2023. The entire study session took about 60 minutes. Participants were compensated with a $50 Amazon gift card.
Each interview began with a brief introduction by the moderator and was divided into three segments: a visualization feedback session with ChatGPT, a similar session with Human expert(s), and an open-ended interview. Six participants initiated the process with the ChatGPT session, while the other six commenced with the Human expert session. After each feedback session, we administered a survey. In five sessions, two visualization experts provided feedback, while in the remaining seven sessions, we had one expert present. Both experts are current professors in the field of visualization.
During the ChatGPT feedback session, participants shared their screens with the moderator and presented their visualizations, accompanied by a brief explanation. The moderator introduced a predefined input format for ChatGPT, employing a role-playing structure (Ihwan, [n. d.]). Participants were guided to furnish details about the visualization’s chart type, textual description, visual encodings, and any related questions. They input this information into ChatGPT while screen-sharing. Following ChatGPT’s response, participants could pose follow-up questions or request clarifications. Afterward, participants received a survey via the Zoom chat to gather feedback on their ChatGPT experience.
During the Human expert session, one or two visualization experts joined the Zoom meeting. Participants shared their screens and presented their visualization-related questions to these experts. The experts answered queries and provided additional feedback or further insights. Subsequently, following this session, the visualization experts left the meeting, and participants were presented with a Zoom survey to gather feedback about their experience.
In the final segment of the study, open-ended interviews were conducted. In these interviews, the moderator inquired about participants’ experiences in both ChatGPT and Human feedback sessions, delving deeper into the rationale behind their survey responses. Furthermore, participants were prompted to share their perspectives on the possible integration of ChatGPT into their data visualization workflow, its constraints, and the future possibilities it might offer.
4.4. Data & Analysis Methods
The final dataset from our user studies comprises participant questions (see Table 2) and responses, survey data, and transcripts from post interviews. Our primary focus for analysis was on the interview transcripts. Initially, two researchers independently examined three interview transcripts, identifying interesting quotes and assigning meaningful themes. Once we established a high degree of consistency between the researchers’ findings, one of them proceeded to review the remaining transcripts. Ultimately, we compiled a categorized list of quotes based on initial higher-level themes (e.g., Where Human experts excel). To further refine our analysis, we revisited these quotes and divided the themes into sub-level categories (e.g., Collaborative and natural conversations).
4.5. Post-Session Interview Analysis Results
| P ID | Chart Type | Initial Questions |
| P10 | parallel coordinate map | Could you suggest an alternative way to visualize this? How can I make this visualization more engaging? |
| P11 | scatterplot | Does it make more sense to compare X or to compare Y? Would there be a better visualization type to show ___? |
| P12 | scatterplot | How can I improve this visualization in order to satisfy the customer? |
| P13 | scatterplot | What is the best way to show ___? what do you think about the current color scheme? |
| P14 | diverging stacked bar chart | How can we ensure the user looks at ___? How can I put these four bars next to each other in one visualization? |
| P15 | bubble cluster | How would I enable more properties to be seen? How would I let the user encode the strength of their preference? |
| P16 | interactive map | Which parts of the data visualization do you think are successful? Are there any areas you believe could be improved for better clarity and impact? |
| P17 | sankey chart | Is there a way to ascertain what visualization is most effective for communicating the data? How could ___ be more clearly displayed? |
| P18 | wav file visualization | Would it be visually overwhelming to use ___? Is this an appropriate visualization for demonstrating ___? |
| P19 | color heatmaps | Would another color map be more suitable? Would you have any suggestions as to how to indicate ___? |
| P20 | beeswarm, line chart | Does the special encoding for ___ make intuitive sense? Is it clear that ___ represents ___? |
| P21 | comparative line graph | How can I decrease the number of colors used in this visualization? Would this data be better represented if I ___? |
4.5.1. Participant Questions and Interaction Dynamics
Four of the participants (P10, P15, P16, and P18) shared interactive visualizations. The other eight participants shared static visualizations. The type of visualization, as well as the questions they initially brought, can be seen in Table 2. To safeguard participant privacy, we refrain from disclosing their visualizations or any elements of their questions that could potentially reveal information about participants.
In the ChatGPT section of the interview, the majority of the participants asked ChatGPT more than one follow-up question. The follow-up questions to ChatGPT tended to be more specific and technical than the follow-up questions asked to the Human experts. For example, P11 asked ChatGPT “What are the pros and cons of having too many bubbles on the scatterplot from having a clear comparison?”
On the other hand, the follow up questions to the Human experts tended to be less structured and more conversational. The participant and expert had a conversation and pulled out different aspects of the question along the way. There was a lot of stopping to clarify what different elements of the visualization meant and how they played into the visualization. As a result, other factors not explicitly mentioned in the question often ended up getting brought up in these sessions with the Human experts.
4.5.2. Post Survey Results
In Figure 4, we present the outcomes of the experience surveys conducted following each feedback session. The overarching consensus among participants was a strong preference for Human experts. This preference was underscored by statistically significant distinctions, as determined through two-tailed Mann-Whitney U tests (1: strongly disagree to 5: strongly agree). Specifically, participants expressed significantly higher levels of satisfaction in interacting with Human experts (, ), perceiving their responses as notably more accurate (, ), helpful (, ), reliable (, ), and adaptable (, ) to their preferences.
Furthermore, respondents indicated that Human experts exhibited a significantly deeper understanding of the context and requirements of their queries. Human experts were also acknowledged for their expertise in data visualization, along with their ability to offer actionable recommendations. It is worth noting that no significant differences were observed in responses to questions related to the clarity and conciseness of explanations, concerns about potential biases, or perceived risks of receiving misleading information.
4.5.3. Where Human Experts Excel
Participants were generally more satisfied with Human experts. They mentioned a variety of reasons for such opinions. The participants highlighted several key themes in their responses, shedding light on the strengths of Human experts in the context of data visualization guidance.
Bespoke and focused feedback of Human experts
First of all, they liked Human experts’ ability to offer tailored recommendations that were closely aligned with the specific visuals in question (P10, P11, P15, P17 - P21). P17 exemplified this sentiment by stating, “[…] recommendations are a lot more tailored and specific in that way, versus just like being tossed a list of potential tools we can look.”. P11 further elaborated on this point by highlighting that Human experts can envision various options and present the ones they deem most valuable. They also noted that Human experts could effectively understand the context of visuals by directly observing them (P11, P17 - P19). P19 expressed this sentiment by stating, “I felt the feedback was obviously more grounded to the visualization at hand. Because again, they could see it.” Similarly, P18 commented, “As he’s looking at the pictures, he knows the information I had access to and how that translates.” Furthermore, participants acknowledged that Human experts excelled in staying focused on the specific problem at hand. For instance, P15 articulated, “[the expert] was able to talk about improvements within scope […] without redesigning it entirely.” Similarly, P10 commented, “The ones that [the expert] gave were within the scope of refining existing visualizations.”
Collaborative and natural conversations
Participants consistently expressed the value of more fluid conversations when interacting with Human experts (P10, P11, P15, P17, P19, P20). They highlighted that Human responses felt more free-flowing, contributing to a sense of genuine interaction (P10, P17). Participants also noted that Human-expert interactions provided a collaborative and interactive experience (P11, P15, P19, P20). For instance, P11 highlighted an efficient turn-taking dynamic by saying, “the ability to ask follow-up questions and clarify I think was easier to do within the Human interaction context.” Others similarly emphasized the convenience of seeking clarifications without the need for precise phrasing or additional context adjustments (P11, P15).
Enriched insights through lived experience
Participants felt that Human experts’ responses were often rooted in their experience and education (P16). They valued the ability of Human experts to explain the underlying rationale, promoting a more actionable approach to implementing the suggestions (P10). Furthermore, participants appreciated the Human experts’ ability to think beyond the obvious; as P13 noted, “I feel like I was definitely more satisfied with the Human experts, […] they just started thinking so many things beyond anything I could have think of.” Moreover, participants conveyed that interactions with Human experts felt personal and supportive. One participant reflected, “they were really walking me through the different things […] help you grow as a professional […] really pushing me.”, while another participant recognized the proactiveness of Human experts by saying “even if I didn’t ask question, they kind of brought it up.”
4.5.4. Where ChatGPT excels
Participants identified the strengths of ChatGPT in contrast to Human experts.
Brainstorming and ideation
ChatGPT’s role as a creative catalyst emerged prominently during participant discussions, underlining its capacity to spark innovative ideas (P11, P15, P18). Participants like P18 found unexpected and impressive suggestions, stating, “It did have some cool ideas, I didn’t expect this from ChatGPT. […] Pretty good.”, while P11 commented, “I think it’s really interesting to try and kind of break outside the box and get more creative… GPT has the potential to disrupt that workflow in a positive way.” Likewise, P15 also acknowledged that the tool may not always provide precise solutions but rather serve as a springboard for creative exploration. P15 contrasts this to Human experts by saying, “[expert] feedback was less brainstorming and more actual critique.”
Broad knowledge base
In a similar vein to its brainstorming capability, participants also highlighted ChatGPT’s capacity to serve as a vast repository of knowledge, offering a wide array of ideas (P10, P11, P15). P10 stated, “it really does kind of feel like a really fancy kind of search engine, where it is kinda, I give it a problem. And it’s like, oh, like, this might be a solution.”, while P15 said, “great place to go to start […] researching or come up with ideas you hadn’t thought of.” P11 similarly emphasized that ChatGPT excels at showcasing a range of data visualization options, making it a valuable resource during the initial stages of projects. P11 elaborated on the tool’s potential, envisioning a scenario where metadata about a dataset could be fed to ChatGPT to receive chart-type recommendations. However, P10 emphasized that a critical eye is needed to discern the relevant and valuable insights from its outputs; “I asked to give 20 ideas and I’m […] experienced enough […] to […] know that, like, 16 of them are just nonsense. The other four things I might not have thought of.”
Time saved via rapid understanding and response
Participants appreciated ChatGPT’s ability to quickly grasp and provide information (P10, P11, P13). P10 said ChatGPT was able to quickly grasp unfamiliar concepts auxiliary to visualization (e.g., explaining Wordle to ChatGPT), while P11 and P13 felt ChatGPT provided faster and more comprehensive lists of options, especially for alternative chart types, when compared to the responses from Human experts. On a related note, P11 said that Human interactions can sometimes involve a lengthier back-and-forth process to arrive at the right question.
Soft attitude and neutral perspective
Other participants noted ChatGPT’s behavioral traits. For example, P19 pointed out that ChatGPT often assumed a gentle and agreeable attitude by acknowledging concerns without offering specific directions, which led to a perception of excessive alignment. P19 also noted that ChatGPT demonstrated a neutral stance, presenting a variety of viewpoints without favoring any particular side. This contrasted with Human experts who might convey their own biases, thereby making ChatGPT’s approach seem more balanced.
4.5.5. Where ChatGPT falls short
Participants also discussed ChatGPT’s several shortcomings in comparison to Human experts.
Lack of depth in responses
Participant feedback indicated a notable lack of depth in ChatGPT’s responses (P13, P15, P19). For instance, P19 expressed that ChatGPT’s advice seemed to lack actionable insights and depth that Human experts possess. This sentiment was echoed by P15, who noted that while ChatGPT’s answers were broad in knowledge, they lacked depth. P15 further elaborated that ChatGPT often suggested radical changes (i.e., diverse ideas), while Human experts provided improvements that were more aligned with refining existing visualizations. P19 similarly commented that ChatGPT’s ideas were not considering the specific visualization. Participants acknowledged this contrast as a trade-off (P15, P17), emphasizing that the desired level of response detail and granularity from either ChatGPT or Human experts depends on user needs.
Generic advice, misalignment, and lack of critical thinking
Participants expressed that ChatGPT’s feedback often felt generic and lacked contextual depth (P12, P16, P20). P20 expressed that ChatGPT’s “knowledge felt […] not necessarily backed up by experience.” P16 and P12 similarly noted that ChatGPT’s responses were not necessarily wrong but were overly generic, failing to consider the actual outcomes of the provided recommendations. P16 further commented that they “felt the areas of improvement were good suggestions. But it didn’t feel like you have to do this.”
Several participants highlighted instances of misalignment and a lack of specificity in ChatGPT’s recommendations (P11, P16, P19). P19 pointed out an example where ChatGPT suggested color maps that might create confusion rather than improve visualizations. This lack of alignment was further emphasized by P11, who noted a communication hurdle hindering the establishment of a common understanding and precise feedback.
Participants highlighted concerns related to critical thinking (P13, P16). P16 noted, “whereas the ChatGPT felt more like it was telling me what I wanted to hear if that makes sense.” P16 further elaborated that ChatGPT’s responses sometimes resembled textbook information rather than offering insightful suggestions on visualization. P13 shared a similar sentiment, stating, “And I feel like the issue with ChatGPT typically is that it only responds to your question. So if you cannot think of the question, it will not give you the answer.”
Efforts required for fluid conversation
Participants highlighted the challenges in having a fluid and interactive conversation with ChatGPT (P10, P13, P16, P17, P20). For instance, P10 explained, “… kind of hoping that one of them would be what I wanted.”. P17 expressed a similar sentiment, stating, “even though it’s still conversational, but it still kind of feels a little bit static sometimes where you’re just typing text into a text box.” P16 further emphasized the difference in interaction dynamics, stating, “I really liked listening to the Human experts go back and forth, and talk about different aspects… whereas ChatGPT just fed that to me, and that was it.” P13 noted that ChatGPT’s responses often required iterations and follow-ups, P10 found that ChatGPT’s insights became more specific and interesting when more constraints were provided.
4.5.6. Opinions on Trust and Reliability with ChatGPT
Participants had varied opinions on the trustworthiness and reliability of ChatGPT’s responses, with some expressing low trust (P10, 18, P20). P20 mentioned the limitations in ChatGPT’s understanding of visuals, which relied heavily on the accuracy of participants’ descriptions. P18 pointed out the potential for fabricated information due to the algorithmic nature of the responses. P10 echoed this sentiment, indicating that ChatGPT often sounded too confident and authoritative, which could mislead users into placing unwarranted trust. On the other hand, P17 said the two are relatively the same, highlighting that the key distinction lies in how information is delivered.
Several other participants pointed to the need for due diligence in evaluating ChatGPT’s suggestions, especially when the user lacks domain expertise (P10, P15). For instance, P10 commented: “If I was in a field that I wasn’t familiar with, I think it’d be really easy to get fooled by it.” P15 drew a parallel with the early inception of Wikipedia, stating, “ it was like early days of Wikipedia, I was taught like, never cite Wikipedia, it could be wrong. And I think the same thing is true of something like a large language model.”
They also made comparisons to Human experts in terms of trust and authority (P13, P16, P18, P21). They explained how the credentials of Human respondents influenced their perception (P18, P21), e.g., “Well, the Human respondent is a professor in visualization.”—P18. Others perceived Human experts as sounding knowledgeable and educational (P13, P16). In contrast to these, participants also shared their perceptions of expertise and self-awareness in AI. P16 said, “Didn’t feel like that the ChatGPT response was incorrect … trusted it, but not nearly as much as the Humans.”, while P20 expressed, “ChatGPT doesn’t know when it’s wrong.”
4.5.7. Limitations and Opportunities with ChatGPT
Participants expressed an optimistic outlook regarding the future potential of ChatGPT, despite acknowledging its current challenges (P10, P12, P15, P16). They discussed aspects they would like to see improved in ChatGPT in the future.
Ability to convey complex visualizations
Participants emphasized the significance of enabling ChatGPT to comprehend and interpret complex visualizations, highlighting the limitations of text-based communication when dealing with intricate design problems (P10, P11, P13, P15 - 21). P15 acknowledged the inherent loss of information when translating between visual and written mediums, while P10 and P11 conveyed the complexity of describing interactive dashboards and multimedia visualizations via text. P10 also mentioned video recordings that could convey the interactivity of designs.
Ability to generate visualizations based on feedback
Several participants pondered the prospect of ChatGPT being able to not only analyze input information but also generate relevant visualizations (P18, P19, P21). For example, P18 proposed the generation or modification of images based on textual descriptions. P19 noted the convenience of sharing underlying code to accompany visual content, emphasizing its value in effectively communicating complex ideas instead of relying solely on text descriptions. Similarly, P21 discussed the concept of inputting data and posing questions to explore the space of chart design. P13 envisioned the integration of ChatGPT within existing tools, enhancing workflow efficiency and facilitating a more cohesive user experience.
Facilitating Fluid and Truthful Conversations
Participants expressed the need for ChatGPT to initiate and support more fluid conversations to enhance its usability and value (P11, P16, P21). P16 highlighted the need for AI to prompt users for further details about the visualization; “I would really like it if it were to ask me questions to kind of gauge my thought process.” Similarly, P11 noted the value of follow-up questions in driving valuable insights, while “it was harder to envision what a follow-up question to that system would be.” On the other hand, P21 raised concerns about the possibility of ChatGPT providing false or hallucinated information.
The indispensable role of Human feedback
The consensus among participants highlighted the enduring value of Human feedback in evaluating and improving visualizations (P11, P16, P21). P16 commented, “I think there’s always, at least in my lifetime, going to be a need for Human interpretation. And the Human experts’ sessions, for me, reinforce that they offer a lot of really good insight.” P11 further expressed that despite ChatGPT’s capabilities, the feedback from actual users remains incomparable, as it encompasses a deeper level of insight into user experience. Additionally, P11 discussed ideas of experimenting with different personas for hypothetical user interactions.
Issues with privacy and sensitive data
Participants in the interviews expressed significant concerns regarding the privacy and handling of sensitive data by ChatGPT (P15, P16, P20, P21). They emphasized that sharing visuals containing private or sensitive information would be uncomfortable and unlikely unless the data was from a public source or there was a guarantee that it would not be shared publicly. P15 shared an illustrative example, stating, “I met a company that I had to sign an NDA with, and we are not allowed to use ChatGPT with anything that might be considered confidential information because that information goes onto OpenAI servers and will live there for eternity.”
4.6. Takeaways
The findings from our study underscore the distinct strengths and limitations of ChatGPT in comparison to Human experts. The clear preference for Human experts among participants highlights their superior ability to provide tailored, context-specific advice and engage in more interactive and collaborative conversations. This preference is rooted in the Human experts’ deep understanding of visualization context and their capacity for nuanced, bespoke feedback.
On the other hand, ChatGPT shows promise as a creative catalyst and a broad knowledge base, offering rapid responses and a diverse array of ideas. However, its limitations are evident in the lack of depth and specificity in responses, challenges in fluid conversation, and concerns about perceived trustworthiness (Yin et al., 2019; Toreini et al., 2020; Drozdal et al., 2020; Yang et al., 2020; Chatzimparmpas et al., 2020) and handling sensitive data (Liu et al., 2021a). Several participants expressed concerns about needing expertise to assess the validity and relevance of potentially inaccurate suggestions from ChatGPT. Nonetheless, they remained optimistic about its future potential, anticipating continued improvements and increasing value.
5. Discussions
5.1. Synthesizing the results of first and second studies
Both studies consistently highlighted similar strengths and weaknesses of ChatGPT compared to Human counterparts. A notable difference was that participants generally showed a stronger preference for Human experts in the second study. This disparity might be attributed to several factors. One reason could be that while any member of the VisGuides community can respond to questions on the forum, the feedback sessions were consistently answered by experienced visualization experts. In generall, the responses from Human experts in these sessions closely paralleled the high-scoring Human responses observed in the Visguides. Moreover, human-to-human interaction might have also played a role. For instance, Human experts ensured that they did not overlook any aspects of participants’ questions, benefiting from real-time conversations. Overall, the second study might demonstrate that human perception of usefulness might extend beyond just response quality.
5.2. Limitations of our studies and future opportunities
While our studies were designed and conducted with considerable care, it is important to acknowledge certain constraints. One notable restriction is the lack of demographic and expertise data for the Human respondents in the VisGuides forum, which poses a challenge in making direct comparisons with the Human experts. Also, we relied on the experiential opinions of participants to assess the feedback sessions, providing more direct and accurate evaluations. Due to the dynamic nature of the feedback sessions, we were not able to apply the static evaluation metrics used in the VisGuides. Despite this limitation, our results were highly consistent.
In this work, we used a fixed version of ChatGPT, a fine-tuned iteration of GPT 3.5. As with many other new technologies, LLMs are continuously advancing. ChatGPT and other LLM-based chatbots are evolving to incorporate more recent data and support new features, including the ability to handle more nuanced language or to understand and generate images, although some of these features are only available to paid users. As a result, responses to the questions in our study will vary among newer LLMs: they can be different, better, or sometimes even worse in certain cases. However, our study results go beyond transient outcomes, as they provide a framework for assessing LLM responses and understanding human-LLM interactions.
Going forward, comprehensive benchmarking of LLMs in data visualization would be beneficial, akin to other domain-specific benchmarks (Fei et al., 2023; Jin et al., 2021; Zhang et al., 2023a). As we anticipate a surge in LLM-driven visualization advancements, ensuring these advancements are grounded in a profound understanding of LLM knowledge capacities would be helpful. Our collection of questions and evaluation metrics can be a stepping stone for future scalable benchmarking for various models beyond ChatGPT. LLM-based evaluation can further address tedious human evaluation issues (Zheng et al., 2023; Liu et al., 2023; Wang et al., 2023; Fu et al., 2023; Kim et al., 2023). This direction in assessing LLM capabilities in data visualization might expand to parallel the evaluation of human abilities in visual perception and cognition (Haehn et al., 2018; Yang et al., 2023).
5.3. Design considerations for LLM-based feedback systems
Our study results indicate the feasibility of developing an advanced design feedback system using LLMs. Drawing from the insights gained, we address several key design considerations for constructing such a system by addressing the strengths and limitations we identified in the study.
Leverage the knowledge base for creative exploration
It was evident that the broad knowledge base of LLMs is advantageous for rapidly exploring diverse ideas or sparking new ideas for participants. However, one issue we observed in the study was that this discovery process was mostly ad-hoc and spontaneous, resulting in a hit-or-miss experience. Future feedback systems have the potential to make this process more systematic and supportive of structured ideation. This could be achieved by providing prompts or guiding questions that steer exploration toward relevant and productive paths. For example, a recent study provided insights into a structured exploration of the design space by generating dimensions to guide response generation (Suh et al., 2023).
Support nuanced understanding of design context
One of the most frequent complaints was the lack of a nuanced understanding of the problem context in the question. LLMs with enhanced image understanding capability will alleviate this issue but will not be enough. Future systems might consider allowing visual prompting in which users can annotate areas of interest. Participants also mentioned generating image responses that can help users envision how to address their design concerns; i.e., these responses could also be annotated to highlight improvements. Moreover, communicating design issues about interactive data visualizations poses interesting research challenges for future systems.
Facilitate fluid and proactive engagements
Our participants favored seamless and dynamic discussions with Human experts. ChatGPT is currently trained to align with users’ intentions (Ouyang et al., 2022), often resulting in passive and affirmative behavior. Future feedback systems can aim to be more proactive, especially for novice users who may struggle to articulate their design concerns, such as by asking clarification questions or suggesting related design questions. Participant comments indicate that this approach may involve a trade-off, such as potentially offering irrelevant feedback when not needed, which can result in a suboptimal user experience. Therefore, it would be beneficial for users to have control over the level of proactiveness.
Improve trustworthiness of design suggestions
Interestingly, although there is no guarantee that Human experts’ responses are more truthful, participants were more concerned about ChatGPT’s credibility. Future systems take inspiration from the behavior of Human experts in our study. Having a fluid back-and-forth conversation to build a common ground was one thing, but they also often accompanied their responses with rationalized explanations and supplemented supporting empirical evidence. In a similar vein, participants were also uneasy about not being able to verify ChatGPT responses especially when they lacked relevant knowledge. Providing application examples, as they do in practice (Choi et al., 2023b; Kang et al., 2018) might alleviate this concern. These abilities might require more than just prompt engineering, such as augmented retrieval (Lewis et al., 2020).
Integrating into a practitioner’s workflow
Several participants mentioned that they already use ChatGPT for tasks related to writing or debugging code (P10, P12, P13, P16, P18). But, data visualization practice rests on two pillars: design and implementation. Our study sheds light on ChatGPT’s capability in terms of design knowledge. As P13 suggests, the integration of a ChatGPT-like assistant within current tools would be highly beneficial. Such a design assistant could suggest appropriate chart types (P11, P21) or provide rationales or critiques for generated visualizations.
5.4. ChatGPT for design knowledge education
Some participants noted ChatGPT’s potential for education (P20, P21), such as assisting with understanding unfamiliar charts or aiding in writing educational blogs on data visualization. A recent paper identifies challenges in visualization education (Bach et al., 2023), highlighting AI as a valuable tool for formal and informal learning contexts. However, concerns include the consistency and quality of AI-generated content, as differing responses to students could be problematic. There’s also the risk of students over-relying on AI instead of engaging critically with visualizations, e.g., through peer feedback. Future AI systems in education may act as personalized tutors, tracking progress, guiding learning goals, and fostering both comprehension and generation of visualizations.
6. Conclusion
In this study, we explored the potential of ChatGPT as a design companion, focusing on its capacity to provide valuable visualization insights to practitioners. Our findings reveal that ChatGPT capitalizes on a vast knowledge repository to generate diverse and imaginative suggestions. While limitations like the contextual depth of its responses and truthful interaction dynamics exist, participants expressed optimism about its future utility. For future work, we plan to investigate ways to build a design feedback system based on the lessons learned. This research direction will involve enhancing chart comprehension and enriching the knowledge base of LLMs with well-founded rationales, including visualization examples and empirical evidence, developing user interfaces and interaction designs tailored to the designer’s workflow, and evaluating the system’s practical applicability.
Acknowledgements.
We would like to acknowledge the support of the National Science Foundation (#2146868).References
- (1)
- dat ([n. d.]) [n. d.]. Data Visualization Society. Accessed: March 28, 2023.
- dvs ([n. d.]) [n. d.]. Data Visualization Society Surveys. https://www.datavisualizationsociety.org/survey-history Accessed on Sep 9, 2023.
- Ahmed et al. (2023) Imtiaz Ahmed, Ayon Roy, Mashrafi Kajol, Uzma Hasan, Partha Protim Datta, and Md Rokonuzzaman Reza. 2023. ChatGPT vs. Bard: a comparative study. Authorea Preprints (2023).
- Bach et al. (2023) Benjamin Bach, Mandy Keck, Fateme Rajabiyazdi, Tatiana Losev, Isabel Meirelles, Jason Dykes, Robert S Laramee, Mashael AlKadi, Christina Stoiber, Samuel Huron, et al. 2023. Challenges and Opportunities in Data Visualization Education: A Call to Action. arXiv preprint arXiv:2308.07703 (2023).
- Baidoo-Anu and Owusu Ansah (2023) David Baidoo-Anu and Leticia Owusu Ansah. 2023. Education in the era of generative artificial intelligence (AI): Understanding the potential benefits of ChatGPT in promoting teaching and learning. Available at SSRN 4337484 (2023).
- Bako et al. (2023) Hannah K. Bako, Xinyi Liu, Leilani Battle, and Zhicheng Liu. 2023. Understanding how Designers Find and Use Data Visualization Examples. IEEE Transactions on Visualization and Computer Graphics 29, 1 (2023), 1048–1058. https://doi.org/10.1109/TVCG.2022.3209490
- Bartolomeo et al. (2023) Sara Di Bartolomeo, Giorgio Severi, Victor Schetinger, and Cody Dunne. 2023. Ask and You Shall Receive (a Graph Drawing): Testing ChatGPT’s Potential to Apply Graph Layout Algorithms. In EuroVis 2023 - Short Papers, Thomas Hoellt, Wolfgang Aigner, and Bei Wang (Eds.). The Eurographics Association. https://doi.org/10.2312/evs.20231047
- Bender et al. (2021) Emily M. Bender, Timnit Gebru, Angelina McMillan-Major, and Shmargaret Shmitchell. 2021. On the Dangers of Stochastic Parrots: Can Language Models Be Too Big?. In Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency (Virtual Event, Canada) (FAccT ’21). Association for Computing Machinery, New York, NY, USA, 610–623. https://doi.org/10.1145/3442188.3445922
- Bommasani et al. (2021) Rishi Bommasani, Drew A. Hudson, Ehsan Adeli, Russ Altman, Simran Arora, Sydney von Arx, Michael S. Bernstein, Jeannette Bohg, Antoine Bosselut, Emma Brunskill, Erik Brynjolfsson, S. Buch, Dallas Card, Rodrigo Castellon, Niladri S. Chatterji, Annie S. Chen, Kathleen A. Creel, Jared Davis, Dora Demszky, Chris Donahue, Moussa Doumbouya, Esin Durmus, Stefano Ermon, John Etchemendy, Kawin Ethayarajh, Li Fei-Fei, Chelsea Finn, Trevor Gale, Lauren E. Gillespie, Karan Goel, Noah D. Goodman, Shelby Grossman, Neel Guha, Tatsunori Hashimoto, Peter Henderson, John Hewitt, Daniel E. Ho, Jenny Hong, Kyle Hsu, Jing Huang, Thomas F. Icard, Saahil Jain, Dan Jurafsky, Pratyusha Kalluri, Siddharth Karamcheti, Geoff Keeling, Fereshte Khani, O. Khattab, Pang Wei Koh, Mark S. Krass, Ranjay Krishna, Rohith Kuditipudi, Ananya Kumar, Faisal Ladhak, Mina Lee, Tony Lee, Jure Leskovec, Isabelle Levent, Xiang Lisa Li, Xuechen Li, Tengyu Ma, Ali Malik, Christopher D. Manning, Suvir P. Mirchandani, Eric Mitchell, Zanele Munyikwa, Suraj Nair, Avanika Narayan, Deepak Narayanan, Benjamin Newman, Allen Nie, Juan Carlos Niebles, Hamed Nilforoshan, J. F. Nyarko, Giray Ogut, Laurel Orr, Isabel Papadimitriou, Joon Sung Park, Chris Piech, Eva Portelance, Christopher Potts, Aditi Raghunathan, Robert Reich, Hongyu Ren, Frieda Rong, Yusuf H. Roohani, Camilo Ruiz, Jack Ryan, Christopher R’e, Dorsa Sadigh, Shiori Sagawa, Keshav Santhanam, Andy Shih, Krishna Parasuram Srinivasan, Alex Tamkin, Rohan Taori, Armin W. Thomas, Florian Tramèr, Rose E. Wang, William Wang, Bohan Wu, Jiajun Wu, Yuhuai Wu, Sang Michael Xie, Michihiro Yasunaga, Jiaxuan You, Matei A. Zaharia, Michael Zhang, Tianyi Zhang, Xikun Zhang, Yuhui Zhang, Lucia Zheng, Kaitlyn Zhou, and Percy Liang. 2021. On the opportunities and risks of foundation models. arXiv preprint arXiv:2108.07258 (2021). https://crfm.stanford.edu/assets/report.pdf
- Borji and Mohammadian (2023) Ali Borji and Mehrdad Mohammadian. 2023. Battle of the Wordsmiths: Comparing ChatGPT, GPT-4, Claude, and Bard. GPT-4, Claude, and Bard (June 12, 2023) (2023).
- Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language Models are Few-Shot Learners. In Advances in Neural Information Processing Systems, H. Larochelle, M. Ranzato, R. Hadsell, M.F. Balcan, and H. Lin (Eds.), Vol. 33. Curran Associates, Inc., 1877–1901. https://proceedings.neurips.cc/paper_files/paper/2020/file/1457c0d6bfcb4967418bfb8ac142f64a-Paper.pdf
- Buhrmester et al. (2011) Michael Buhrmester, Tracy Kwang, and Samuel D. Gosling. 2011. Amazon’s Mechanical Turk: A New Source of Inexpensive, Yet High-Quality, Data? Perspectives on Psychological Science 6, 1 (2011), 3–5. https://doi.org/10.1177/1745691610393980 PMID: 26162106.
- Chalkidis et al. (2019) Ilias Chalkidis, Ion Androutsopoulos, and Nikolaos Aletras. 2019. Neural Legal Judgment Prediction in English. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics. Association for Computational Linguistics, Florence, Italy, 4317–4323. https://doi.org/10.18653/v1/P19-1424
- Chang et al. (2023) Yupeng Chang, Xu Wang, Jindong Wang, Yuan Wu, Kaijie Zhu, Hao Chen, Linyi Yang, Xiaoyuan Yi, Cunxiang Wang, Yidong Wang, et al. 2023. A survey on evaluation of large language models. arXiv preprint arXiv:2307.03109 (2023).
- Chatzimparmpas et al. (2020) Angelos Chatzimparmpas, Rafael Messias Martins, Ilir Jusufi, Kostiantyn Kucher, Fabrice Rossi, and Andreas Kerren. 2020. The state of the art in enhancing trust in machine learning models with the use of visualizations. In Computer Graphics Forum, Vol. 39. Wiley Online Library, 713–756.
- Choi et al. (2023b) Jinhan Choi, Changhoon Oh, Yea-Seul Kim, and Nam Wook Kim. 2023b. VisLab: Enabling Visualization Designers to Gather Empirically Informed Design Feedback. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 813, 18 pages. https://doi.org/10.1145/3544548.3581132
- Choi et al. (2021) Jinhan Choi, Changhoon Oh, Bongwon Suh, and Nam Wook Kim. 2021. Toward a Unified Framework for Visualization Design Guidelines. In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI EA ’21). Association for Computing Machinery, New York, NY, USA, Article 240, 7 pages. https://doi.org/10.1145/3411763.3451702
- Choi et al. (2023a) Jonathan H Choi, Kristin E Hickman, Amy Monahan, and Daniel Schwarcz. 2023a. Chatgpt goes to law school. Available at SSRN (2023).
- Devlin et al. (2018) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2018. Bert: Pre-training of deep bidirectional transformers for language understanding. arXiv preprint arXiv:1810.04805 (2018).
- Dibia (2023) Victor Dibia. 2023. LIDA: A Tool for Automatic Generation of Grammar-Agnostic Visualizations and Infographics using Large Language Models. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 3: System Demonstrations). Association for Computational Linguistics, Toronto, Canada, 113–126. https://doi.org/10.18653/v1/2023.acl-demo.11
- Diehl et al. (2018) Alexandra Diehl, Alfie Abdul-Rahman, Mennatallah El-Assady, Benjamin Bach, Daniel Keim, and Min Chen. 2018. VisGuides: A Forum for Discussing Visualization Guidelines. In EuroVis 2018 - Short Papers, Jimmy Johansson, Filip Sadlo, and Tobias Schreck (Eds.). The Eurographics Association. https://doi.org/10.2312/eurovisshort.20181079
- Diehl et al. (2021) Alexandra Diehl, Elif E. Firat, Thomas Torsney-Weir, Alfie Abdul-Rahman, Benjamin Bach, Robert Laramee, Renato Pajarola, and Min Chen. 2021. VisGuided: A Community-driven Approach for Education in Visualization. In Eurographics 2021 - Education Papers, Beatriz Sousa Santos and Gitta Domik (Eds.). The Eurographics Association. https://doi.org/10.2312/eged.20211003
- Drozdal et al. (2020) Jaimie Drozdal, Justin Weisz, Dakuo Wang, Gaurav Dass, Bingsheng Yao, Changruo Zhao, Michael Muller, Lin Ju, and Hui Su. 2020. Trust in AutoML: exploring information needs for establishing trust in automated machine learning systems. In Proceedings of the 25th international conference on intelligent user interfaces. 297–307.
- Esteves and Neves (2022) Salomé Esteves and Marco Neves. 2022. “I learned it on the job” Becoming a Data Visualization professional in news media. Information Design Journal 27, 3 (2022), 309–319. https://doi.org/10.1075/idj.22004.est
- Fei et al. (2023) Zhiwei Fei, Xiaoyu Shen, Dawei Zhu, Fengzhe Zhou, Zhuo Han, Songyang Zhang, Kai Chen, Zongwen Shen, and Jidong Ge. 2023. LawBench: Benchmarking Legal Knowledge of Large Language Models. arXiv preprint arXiv:2309.16289 (2023).
- Frosolini et al. (2023) Andrea Frosolini, Paolo Gennaro, Flavia Cascino, and Guido Gabriele. 2023. In Reference to “Role of Chat GPT in Public Health”, to Highlight the AI’s Incorrect Reference Generation. Annals of Biomedical Engineering (2023), 1–3.
- Fu et al. (2023) Jinlan Fu, See-Kiong Ng, Zhengbao Jiang, and Pengfei Liu. 2023. Gptscore: Evaluate as you desire. arXiv preprint arXiv:2302.04166 (2023).
- Gehman et al. (2020) Samuel Gehman, Suchin Gururangan, Maarten Sap, Yejin Choi, and Noah A Smith. 2020. RealToxicityPrompts: Evaluating Neural Toxic Degeneration in Language Models. In Findings of the Association for Computational Linguistics: EMNLP 2020. Association for Computational Linguistics, 3356–3369. https://doi.org/10.18653/v1/2020.findings-emnlp.301
- Giannakopoulos et al. (2023) Kostis Giannakopoulos, Argyro Kavadella, Anas Aaqel Salim, Vassilis Stamatopoulos, and Eleftherios G Kaklamanos. 2023. Evaluation of the Performance of Generative AI Large Language Models ChatGPT, Google Bard, and Microsoft Bing Chat in Supporting Evidence-Based Dentistry: Comparative Mixed Methods Study. Journal of medical Internet research 25 (2023), e51580.
- Gilson et al. (2023) Aidan Gilson, Conrad W Safranek, Thomas Huang, Vimig Socrates, Ling Chi, Richard Andrew Taylor, David Chartash, et al. 2023. How does ChatGPT perform on the United States medical licensing examination? The implications of large language models for medical education and knowledge assessment. JMIR Medical Education 9, 1 (2023), e45312. https://doi.org/10.1101/2022.12.23.22283901
- Haehn et al. (2018) Daniel Haehn, James Tompkin, and Hanspeter Pfister. 2018. Evaluating ‘graphical perception’with CNNs. IEEE transactions on visualization and computer graphics 25, 1 (2018), 641–650.
- Haug and Mädche (2021) Saskia Haug and Alexander Mädche. 2021. Crowd-Feedback in Information Systems Development: A State-of-the-Art Review. In ICIS 2021 Proceedings. Association for Information Systems (AIS). https://doi.org/10.5445/IR/1000139669
- Hutchinson et al. (2023) Maeve Hutchinson, Aidan Slingsby, Radu Jianu, and Pranava Madhyastha. 2023. Towards Visualisation Specifications from Multilingual Natural Language Queries using Large Language Models. (2023). https://doi.org/10.2312/evp.20231072
- Ihwan ([n. d.]) Aris Ihwan. [n. d.]. Role Prompting. https://www.linkedin.com/pulse/role-prompting-aris-ihwan/. Accessed on Sep 8, 2023.
- Jin et al. (2021) Di Jin, Eileen Pan, Nassim Oufattole, Wei-Hung Weng, Hanyi Fang, and Peter Szolovits. 2021. What disease does this patient have? a large-scale open domain question answering dataset from medical exams. Applied Sciences 11, 14 (2021), 6421.
- Kang et al. (2018) Hyeonsu B. Kang, Gabriel Amoako, Neil Sengupta, and Steven P. Dow. 2018. Paragon: An Online Gallery for Enhancing Design Feedback with Visual Examples. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (Montreal QC, Canada) (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–13. https://doi.org/10.1145/3173574.3174180
- Kasneci et al. (2023) Enkelejda Kasneci, Kathrin Sessler, Stefan Küchemann, Maria Bannert, Daryna Dementieva, Frank Fischer, Urs Gasser, Georg Groh, Stephan Günnemann, Eyke Hüllermeier, Stephan Krusche, Gitta Kutyniok, Tilman Michaeli, Claudia Nerdel, Jürgen Pfeffer, Oleksandra Poquet, Michael Sailer, Albrecht Schmidt, Tina Seidel, Matthias Stadler, Jochen Weller, Jochen Kuhn, and Gjergji Kasneci. 2023. ChatGPT for good? On opportunities and challenges of large language models for education. Learning and Individual Differences 103 (2023), 102274. https://doi.org/10.1016/j.lindif.2023.102274
- Kim et al. (2023) Seungone Kim, Jamin Shin, Yejin Cho, Joel Jang, Shayne Longpre, Hwaran Lee, Sangdoo Yun, Seongjin Shin, Sungdong Kim, James Thorne, et al. 2023. Prometheus: Inducing fine-grained evaluation capability in language models. arXiv preprint arXiv:2310.08491 (2023).
- Krause et al. (2017) Markus Krause, Tom Garncarz, JiaoJiao Song, Elizabeth M. Gerber, Brian P. Bailey, and Steven P. Dow. 2017. Critique Style Guide: Improving Crowdsourced Design Feedback with a Natural Language Model. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI ’17). Association for Computing Machinery, New York, NY, USA, 4627–4639. https://doi.org/10.1145/3025453.3025883
- Lahat et al. (2023) Adi Lahat, Eyal Shachar, Benjamin Avidan, Zina Shatz, Benjamin S. Glicksberg, and Eyal Klang. 2023. Evaluating the use of large language model in identifying top research questions in gastroenterology. Scientific Reports 13, 1 (March 2023), 4164. https://doi.org/10.1038/s41598-023-31412-2
- Lee et al. (2023) Peter Lee, Sebastien Bubeck, and Joseph Petro. 2023. Benefits, limits, and risks of GPT-4 as an AI chatbot for medicine. New England Journal of Medicine 388, 13 (2023), 1233–1239. https://doi.org/10.1056/NEJMsr2214184
- Lewis et al. (2020) Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen-tau Yih, Tim Rocktäschel, et al. 2020. Retrieval-augmented generation for knowledge-intensive nlp tasks. Advances in Neural Information Processing Systems 33 (2020), 9459–9474.
- Liang et al. (2022) Percy Liang, Rishi Bommasani, Tony Lee, Dimitris Tsipras, Dilara Soylu, Michihiro Yasunaga, Yian Zhang, Deepak Narayanan, Yuhuai Wu, Ananya Kumar, et al. 2022. Holistic evaluation of language models. arXiv preprint arXiv:2211.09110 (2022).
- Lim et al. (2023) Zhi Wei Lim, Krithi Pushpanathan, Samantha Min Er Yew, Yien Lai, Chen-Hsin Sun, Janice Sing Harn Lam, David Ziyou Chen, Jocelyn Hui Lin Goh, Marcus Chun Jin Tan, Bin Sheng, et al. 2023. Benchmarking large language models’ performances for myopia care: a comparative analysis of ChatGPT-3.5, ChatGPT-4.0, and Google Bard. EBioMedicine 95 (2023).
- Liu et al. (2021a) Bo Liu, Ming Ding, Sina Shaham, Wenny Rahayu, Farhad Farokhi, and Zihuai Lin. 2021a. When machine learning meets privacy: A survey and outlook. ACM Computing Surveys (CSUR) 54, 2 (2021), 1–36.
- Liu et al. (2021b) Pengfei Liu, Weizhe Yuan, Jinlan Fu, Zhengbao Jiang, Hiroaki Hayashi, and Graham Neubig. 2021b. Pre-train, prompt, and predict: A systematic survey of prompting methods in natural language processing. arXiv 2021. arXiv preprint arXiv:2107.13586 (2021).
- Liu et al. (2023) Yang Liu, Dan Iter, Yichong Xu, Shuohang Wang, Ruochen Xu, and Chenguang Zhu. 2023. Gpteval: Nlg evaluation using gpt-4 with better human alignment. arXiv preprint arXiv:2303.16634 (2023).
- Luther et al. (2014) Kurt Luther, Amy Pavel, Wei Wu, Jari-lee Tolentino, Maneesh Agrawala, Björn Hartmann, and Steven P. Dow. 2014. CrowdCrit: Crowdsourcing and Aggregating Visual Design Critique. In Proceedings of the Companion Publication of the 17th ACM Conference on Computer Supported Cooperative Work & Social Computing (Baltimore, Maryland, USA) (CSCW Companion ’14). Association for Computing Machinery, New York, NY, USA, 21–24. https://doi.org/10.1145/2556420.2556788
- Luther et al. (2015) Kurt Luther, Jari-Lee Tolentino, Wei Wu, Amy Pavel, Brian P. Bailey, Maneesh Agrawala, Björn Hartmann, and Steven P. Dow. 2015. Structuring, Aggregating, and Evaluating Crowdsourced Design Critique. In Proceedings of the 18th ACM Conference on Computer Supported Cooperative Work & Social Computing (Vancouver, BC, Canada) (CSCW ’15). Association for Computing Machinery, New York, NY, USA, 473–485. https://doi.org/10.1145/2675133.2675283
- Mollman (2023) Steve Mollman. 2023. ChatGPT passed a Wharton MBA exam and it’s still in its infancy. One professor is sounding the alarm. https://fortune.com/2023/01/21/chatgpt-passed-wharton-mba-exam-one-professor-is-sounding-alarm-artificial-intelligence/. Accessed on Sep 7, 2023.
- Nadeem et al. (2021) Moin Nadeem, Anna Bethke, and Siva Reddy. 2021. StereoSet: Measuring stereotypical bias in pretrained language models. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers). Association for Computational Linguistics, 5356–5371. https://doi.org/10.18653/v1/2021.acl-long.416
- Okerlund et al. (2022) Johanna Okerlund, Evan Klasky, Aditya Middha, Sujin Kim, Hannah Rosenfeld, Molly Kleinman, and Shobita Parthasarathy. 2022. What’s in the Chatterbox? Large Language Models, Why They Matter, and What We Should Do About Them.
- OpenAI ([n. d.]) OpenAI. [n. d.]. ChatGPT. https://openai.com/chatgpt. Accessed on Sep 7, 2023.
- OpenAI (2023) R OpenAI. 2023. GPT-4 technical report. arXiv (2023), 2303–08774.
- Ouyang et al. (2022) Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, et al. 2022. Training language models to follow instructions with human feedback. Advances in Neural Information Processing Systems 35 (2022), 27730–27744.
- Paik (2023) Soonk Paik. 2023. How I Created a Data Visualization With Zero Coding Skills, Thanks to ChatGPT. Nightingale magazine by the Data Visualization Society (4 April 2023). https://nightingaledvs.com/data-visualization-using-chatgpt-to-code/
- Pan (2023) Jie Pan. 2023. Large language model for molecular chemistry. Nature Computational Science 3, 1 (2023), 5–5. https://doi.org/10.1038/s43588-023-00399-1
- Parsons (2022) P. Parsons. 2022. Understanding Data Visualization Design Practice. IEEE Transactions on Visualization; Computer Graphics 28, 01 (jan 2022), 665–675. https://doi.org/10.1109/TVCG.2021.3114959
- Parsons et al. (2020) Paul Parsons, Colin M. Gray, Ali Baigelenov, and Ian Carr. 2020. Design Judgment in Data Visualization Practice. In 2020 IEEE Visualization Conference (VIS). 176–180. https://doi.org/10.1109/VIS47514.2020.00042
- Parsons and Shukla (2020) Paul Parsons and Prakash Shukla. 2020. Data Visualization Practitioners’ Perspectives on Chartjunk. In 2020 IEEE Visualization Conference (VIS). 211–215. https://doi.org/10.1109/VIS47514.2020.00049
- Qin et al. (2023) Chengwei Qin, Aston Zhang, Zhuosheng Zhang, Jiaao Chen, Michihiro Yasunaga, and Diyi Yang. 2023. Is ChatGPT a general-purpose natural language processing task solver? arXiv preprint arXiv:2302.06476 abs/2302.06476 (2023).
- Radford et al. (2018) Alec Radford, Karthik Narasimhan, Tim Salimans, Ilya Sutskever, et al. 2018. Improving language understanding by generative pre-training. (2018).
- Radford et al. (2019) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya Sutskever, et al. 2019. Language models are unsupervised multitask learners. OpenAI blog 1, 8 (2019), 9.
- Schetinger et al. (2023) V. Schetinger, S. Di Bartolomeo, M. El-Assady, A. McNutt, M. Miller, J. P. A. Passos, and J. L. Adams. 2023. Doom or Deliciousness: Challenges and Opportunities for Visualization in the Age of Generative Models. Computer Graphics Forum 42, 3 (2023), 423–435. https://doi.org/10.1111/cgf.14841 arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1111/cgf.14841
- Shin et al. (2023) Sungbok Shin, Sanghyun Hong, and Niklas Elmqvist. 2023. Perceptual Pat: A Virtual Human Visual System for Iterative Visualization Design. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–17.
- Suh et al. (2023) Sangho Suh, Meng Chen, Bryan Min, Toby Jia-Jun Li, and Haijun Xia. 2023. Structured Generation and Exploration of Design Space with Large Language Models for Human-AI Co-Creation. arXiv preprint arXiv:2310.12953 (2023).
- Tan and Celis (2019) Yi Chern Tan and L. Elisa Celis. 2019. Assessing Social and Intersectional Biases in Contextualized Word Representations. Curran Associates Inc., Red Hook, NY, USA.
- Tlili et al. (2023) Ahmed Tlili, Boulus Shehata, Michael Agyemang Adarkwah, Aras Bozkurt, Daniel T. Hickey, Ronghuai Huang, and Brighter Agyemang. 2023. What if the devil is my guardian angel: ChatGPT as a case study of using chatbots in education. Smart Learning Environments 10, 1 (2023), 15. https://doi.org/10.1186/s40561-023-00237-x
- Toreini et al. (2020) Ehsan Toreini, Mhairi Aitken, Kovila Coopamootoo, Karen Elliott, Carlos Gonzalez Zelaya, and Aad Van Moorsel. 2020. The relationship between trust in AI and trustworthy machine learning technologies. In Proceedings of the 2020 conference on fairness, accountability, and transparency. 272–283.
- Touvron et al. (2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023. Llama: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971 (2023).
- Vaithilingam et al. (2022) Priyan Vaithilingam, Tianyi Zhang, and Elena L. Glassman. 2022. Expectation vs. Experience: Evaluating the Usability of Code Generation Tools Powered by Large Language Models. In Extended Abstracts of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI EA ’22). Association for Computing Machinery, New York, NY, USA, Article 332, 7 pages. https://doi.org/10.1145/3491101.3519665
- Wang et al. (2023) Jiaan Wang, Yunlong Liang, Fandong Meng, Haoxiang Shi, Zhixu Li, Jinan Xu, Jianfeng Qu, and Jie Zhou. 2023. Is chatgpt a good nlg evaluator? a preliminary study. arXiv preprint arXiv:2303.04048 (2023).
- Wang et al. (2022) Xuezhi Wang, Jason Wei, Dale Schuurmans, Quoc Le, Ed Chi, Sharan Narang, Aakanksha Chowdhery, and Denny Zhou. 2022. Self-consistency improves chain of thought reasoning in language models. arXiv preprint arXiv:2203.11171 (2022).
- Wei et al. (2022a) Jason Wei, Yi Tay, Rishi Bommasani, Colin Raffel, Barret Zoph, Sebastian Borgeaud, Dani Yogatama, Maarten Bosma, Denny Zhou, Donald Metzler, Ed H. Chi, Tatsunori Hashimoto, Oriol Vinyals, Percy Liang, Jeff Dean, and William Fedus. 2022a. Emergent abilities of large language models. Transactions on Machine Learning Research (2022).
- Wei et al. (2022b) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022b. Chain-of-thought prompting elicits reasoning in large language models. Advances in Neural Information Processing Systems 35 (2022), 24824–24837.
- Wu et al. (2023) Tianyu Wu, Shizhu He, Jingping Liu, Siqi Sun, Kang Liu, Qing-Long Han, and Yang Tang. 2023. A brief overview of ChatGPT: The history, status quo and potential future development. IEEE/CAA Journal of Automatica Sinica 10, 5 (2023), 1122–1136.
- Xu et al. (2014) Anbang Xu, Shih-Wen Huang, and Brian Bailey. 2014. Voyant: Generating Structured Feedback on Visual Designs Using a Crowd of Non-Experts. In Proceedings of the 17th ACM Conference on Computer Supported Cooperative Work & Social Computing (Baltimore, Maryland, USA) (CSCW ’14). Association for Computing Machinery, New York, NY, USA, 1433–1444. https://doi.org/10.1145/2531602.2531604
- Yang et al. (2020) Fumeng Yang, Zhuanyi Huang, Jean Scholtz, and Dustin L Arendt. 2020. How do visual explanations foster end users’ appropriate trust in machine learning?. In Proceedings of the 25th international conference on intelligent user interfaces. 189–201.
- Yang et al. (2023) Fumeng Yang, Yuxin Ma, Lane Harrison, James Tompkin, and David H. Laidlaw. 2023. How Can Deep Neural Networks Aid Visualization Perception Research? Three Studies on Correlation Judgments in Scatterplots. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Hamburg¡/city¿, ¡country¿Germany¡/country¿, ¡/conf-loc¿) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 822, 17 pages. https://doi.org/10.1145/3544548.3581111
- Yen et al. (2016) Yu-Chun Yen, Steven P Dow, Elizabeth Gerber, and Brian P Bailey. 2016. Social network, web forum, or task market? Comparing different crowd genres for design feedback exchange. In Proceedings of the 2016 ACM Conference on Designing Interactive Systems. 773–784.
- Yin et al. (2019) Ming Yin, Jennifer Wortman Vaughan, and Hanna Wallach. 2019. Understanding the effect of accuracy on trust in machine learning models. In Proceedings of the 2019 chi conference on human factors in computing systems. 1–12.
- Yuan et al. (2016) Alvin Yuan, Kurt Luther, Markus Krause, Sophie Isabel Vennix, Steven P Dow, and Bjorn Hartmann. 2016. Almost an Expert: The Effects of Rubrics and Expertise on Perceived Value of Crowdsourced Design Critiques. In Proceedings of the 19th ACM Conference on Computer-Supported Cooperative Work & Social Computing (San Francisco, California, USA) (CSCW ’16). Association for Computing Machinery, New York, NY, USA, 1005–1017. https://doi.org/10.1145/2818048.2819953
- Zhang et al. (2023a) Tianyi Zhang, Faisal Ladhak, Esin Durmus, Percy Liang, Kathleen McKeown, and Tatsunori B Hashimoto. 2023a. Benchmarking large language models for news summarization. arXiv preprint arXiv:2301.13848 (2023).
- Zhang et al. (2023b) Zhuosheng Zhang, Aston Zhang, Mu Li, Hai Zhao, George Karypis, and Alex Smola. 2023b. Multimodal chain-of-thought reasoning in language models. arXiv preprint arXiv:2302.00923 (2023).
- Zhao et al. (2023) Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, Yifan Du, Chen Yang, Yushuo Chen, Zhipeng Chen, Jinhao Jiang, Ruiyang Ren, Yifan Li, Xinyu Tang, Zikang Liu, Peiyu Liu, Jian-Yun Nie, and Ji-Rong Wen. 2023. A survey of large language models. arXiv preprint arXiv:2303.18223 (2023).
- Zheng et al. (2023) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric Xing, et al. 2023. Judging LLM-as-a-judge with MT-Bench and Chatbot Arena. arXiv preprint arXiv:2306.05685 (2023).