### v4

# How Good is ChatGPT in Giving Advice on Your Visualization Design?

Abstract.
Data visualization creators often lack formal training, resulting in a knowledge gap in design practice. Large-language models such as ChatGPT, with their vast internet-scale training data, offer transformative potential to address this gap. In this study, we used both qualitative and quantitative methods to investigate how well ChatGPT can address visualization design questions. First, we quantitatively compared the ChatGPT-generated responses with anonymous online Human replies to data visualization questions on the VisGuides user forum. Next, we conducted a qualitative user study examining the reactions and attitudes of practitioners toward ChatGPT as a visualization design assistant. Participants were asked to bring their visualizations and design questions and received feedback from both Human experts and ChatGPT in randomized order. Our findings from both studies underscore ChatGPT’s strengths—particularly its ability to rapidly generate diverse design options—while also highlighting areas for improvement, such as nuanced contextual understanding and fluid interaction dynamics beyond the chat interface. Drawing on these insights, we discuss design considerations for future LLM-based design feedback systems.
data visualization, design feedback, ChatGPT, LLM, AI
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Human-centered computing Empirical studies in visualizationccs: Human-centered computing User studies
1. Introduction
Visualizations are ubiquitous and widely employed by practitioners across various disciplines. However, many of those creating visualizations commonly lack formal training in visualization design and instead acquire the necessary skills on the go, leaning from examples from the internet, online blogs, and other publicly available resources (Esteves and Neves, 2022; dvs, [n. d.]). These people may find themselves in situations where they must navigate intricate design choices (Choi et al., 2023b) and consequently revert to their instincts, drawing from the experiences and observations they have accumulated along the way (Choi et al., 2023b; Bako et al., 2023). Alternatively, they seek feedback from colleagues who can help provide novel perspectives, validate design choices, and challenge assumptions (Choi et al., 2023b; Luther et al., 2015). Nevertheless, not all practitioners have the privilege of accessing such feedback from experienced colleagues.
This paper delves into ways to address this knowledge gap in practical settings. We examine the capabilities of large language models (LLMs) equipped with extensive internet-scale training data, which holds transformative potential (Okerlund et al., 2022). LLM-based chatbots, exemplified by ChatGPT (OpenAI, [n. d.]), can offer tailored suggestions by analyzing a practitioner’s design choices and considering established best practices. This emerging advancement can assist casual and novice practitioners, who may have limited knowledge and resources in data visualization, by providing human-like guidance throughout their design journey.
However, open questions remain regarding the extent of LLMs’ current knowledge of visualization best practices, the quality and practical applicability of their suggestions, and how these suggestions differ from those provided by human counterparts. Addressing these open questions, we have formulated two following research questions to guide our investigation:
  • RQ1 (Quantitative evaluation): How well does ChatGPT respond to visualization design questions posed in online communities compared to practitioners in the wild?
  • RQ2 (Qualitative experience): How do visualization practitioners perceive ChatGPT’s design feedback compared to those offered by seasoned professionals?
Figure 1. Methodology Overview: The methodology comprises two key phases. In the first phase, questions answered within a forum space by Human respondents are explored and then presented to ChatGPT. The second phase involves a feedback session in which users solicit visualization design feedback from both ChatGPT and Human experts.
We employ a mixed-methods approach (Fig. 1) to tackle these questions, capturing both breadth (RQ1) and depth (RQ2). RQ1 focuses on a more systematic evaluation using a broader and more realistic knowledge baseline, while RQ2 centers on the qualitative experience of receiving feedback from ChatGPT, assessed against the high standards of domain experts. In our comparisons, seasoned professionals represent trusted colleagues who bring domain expertise and rigor, whereas online practitioners provide a practical, real-world perspective, recognizing that not everyone has access to expert advice (Choi et al., 2023b). We selected ChatGPT as a representative proxy for LLMs due to its widespread use and generally superior performance compared to other LLMs (Borji and Mohammadian, 2023; Ahmed et al., 2023; Giannakopoulos et al., 2023; Lim et al., 2023).
For the first question, we investigated posts and discussions at the VisGuides forum (Diehl et al., 2018), which captures public questions and responses about visualization design. We fed the questions to ChatGPT and compared its responses to the Human responses on VisGuides across six key metrics: coverage, topicality, breadth, clarity, depth, and actionability. We found that ChatGPT’s response is comparable to, and often better than, Human responses, especially in providing clear and comprehensive answers that cover all questions asked. In contrast, human responses typically offered specific suggestions, like citing references beyond basic improvements, but showed greater variability due to multiple respondents with diverse backgrounds involved. Given our analysis pertains to specific versions of ChatGPT, we discuss how we can devise an evaluation pipeline that uses LLMs for automatic response scoring.
For the second question, we conducted a qualitative user study in which we invited practitioners to participate in a comparative evaluation of Human expert feedback and ChatGPT-generated feedback (GPT-3.5 Turbo). For this study, participants were invited to bring their own visualization designs and projects to our sessions for questions and feedback. Most study participants preferred discussing with Human experts, valuing their ability to engage in fluid conversations and provide tailored recommendations. However, participants also appreciated the broad knowledge of ChatGPT, particularly its ability for quick brainstorming and ideation from a neutral standpoint, while acknowledging its limitations in deeper contextual understanding and its lack of critical feedback.
Both studies consistently shed light on the strengths and weaknesses of ChatGPT and Humans. ChatGPT excelled at providing broad suggestions, whereas Humans demonstrated a better grasp of nuanced details for more in-depth ideas. A stronger preference for Humans in the feedback study was also linked to the more consistent opinions and fluid interaction dynamics of Human experts. These findings also point to areas for future enhancement of LLMs, such as supporting advanced prompting with better visual perception and fostering proactive and critical discussion capabilities to more effectively handle the intricate contexts of design challenges. We distill these lessons into design considerations for future design feedback systems using LLMs. Furthermore, we discuss the potential role of an LLM-based design knowledge companion in visualization education.
Our work represents a meaningful step forward for the field amidst the rise of LLM-driven visualization innovations. The contributions of this work lies in the following:
  • Evaluation framework for assessing the capabilities of LLMs in providing data visualization design assistance, incorporating both qualitative and quantitative metrics.
  • Key insights into the role of LLM-based chatbots, exemplified by ChatGPT, in visualization practice, highlighting their strengths and limitations compared to Human counterparts.
  • Design considerations for improving the user interface of LLM-based chatbots as collaborative design companions, fostering more effective human-AI interaction.
2. Related Work
2.1. Data Visualization Design Practice
Data visualization has gone mainstream, frequently used to explore and communicate data in a wide variety of industry sectors. As a result, data visualization practitioners have diverse professions, such as managers, scientists, journalists, and analysts, not necessarily having a specific job title as “visualization designer” or “visualization developer.” (Esteves and Neves, 2022) These professionals do not typically have formal training or education in data visualization, while mostly learning necessary visualization skills on the go (Esteves and Neves, 2022; dvs, [n. d.]). While decades of empirical research in data visualization have produced fruitful knowledge in effective visualization design, it is unclear how well it is accessible to practice (Choi et al., 2021; Diehl et al., 2018). In fact, it has been reported that many practitioners are not aware of theories and principles from research (Parsons, 2022).
As with any creative design profession, visualization practitioners run into challenging situations where they must make decisions among conflicting design alternatives (Choi et al., 2023b). Such decisions might include selecting chart types, making competing decisions between aesthetics and functions, picking appropriate color scales, and addressing conflicts with user and business stakeholder needs (Choi et al., 2023b). Processes by practitioners to address such design issues are more nuanced and multi-faceted than theoretical process models proposed in the research field (Parsons, 2022). They depend on a kind of situated planning that often relies on forms of experimentation and responding to what is happening in the moment (Parsons, 2022; Parsons et al., 2020). They balance readability and engagement and also consider contextual factors such as whether their visualizations are business-oriented dashboards or public-facing graphics (Parsons and Shukla, 2020).
Practitioners frequently resort to their intuition or gut instinct while also looking for examples for inspiration (Choi et al., 2023b; Parsons, 2022). Moreover, they also seek feedback from their colleagues to improve their data visualizations (Choi et al., 2023b; Esteves and Neves, 2022). Such feedback is essential to assess a design and generate revision ideas in the design process, although the fear of criticism and non-anonymity can make people uncomfortable receiving feedback from colleagues. Moreover, self-employed people might struggle to find such colleagues who can provide valuable feedback (Chalkidis et al., 2019)
Previous research has utilized online crowdsourcing to streamline the feedback collection process, providing rapid access to large, diverse participant pools and enabling faster feedback cycles at reduced costs (Buhrmester et al., 2011; Haug and Mädche, 2021; Yen et al., 2016; Luther et al., 2015). To address the challenge of insufficient expertise among crowd workers, researchers have explored various techniques to enhance feedback quality, including micro-tasking (Xu et al., 2014; Luther et al., 2014), implementing predefined rubrics (Yuan et al., 2016), providing illustrative examples (Kang et al., 2018), and establishing feedback guidelines (Krause et al., 2017). While these efforts have primarily focused on feedback for graphic design projects, recent studies in data visualization have investigated automated methods for providing perceptual feedback through image processing techniques (Shin et al., 2023) and developing structured methods for collecting empirically-driven quantitative feedback (Choi et al., 2023b).
In this work, we examine how an LLM-based chatbot, trained with extensive knowledge in the wild, can serve as a design companion to provide knowledge and feedback to varying backgrounds of practitioners. Can LLMs replace the feedback gathering process? We aim to bridge the gap between LLMs’ theoretical potential and their practical application in data visualization, a critical endeavor in the evolving landscape of artificial intelligence within the field.
2.2. Evaluating LLMs’ Knowledge Capacity
LLMs, such as BERT (Devlin et al., 2018), GPT (Brown et al., 2020; OpenAI, 2023), and Llama (Touvron et al., 2023), have rapidly gained popularity and adoption across both industry and academia (Bommasani et al., 2021; Zhao et al., 2023). Leveraging massive amounts of text data and computational power, these foundation models have achieved impressive performance on a wide variety of tasks, from generating human-like text to high-quality images (Radford et al., 2019, 2018; Wei et al., 2022a). However, the scale and complexity of these models make it difficult to fully understand the scope and limitations of their language capabilities (Bender et al., 2021).
A number of benchmark tasks have been proposed to assess the general knowledge and reasoning capabilities of language models in a more holistic manner (Liang et al., 2022; Chang et al., 2023). These include tasks that require models to answer broad-coverage factoid questions based on pre-existing knowledge (Liang et al., 2022). Other benchmarks evaluate numeric, analogical, and logical reasoning skills across different modalities (Qin et al., 2023; Zhang et al., 2023), while some others examine the ability of LLMs to generalize knowledge to new concepts (Wei et al., 2022a). Assessing models on such diverse reasoning skills can provide deeper insight into their adaptability, knowledge gaps, and limitations. In addition to assessing such capabilities, an important area of research has focused on evaluating potential harms and biases in large language models. Some studies have performed targeted adversarial evaluations to expose problematic biases encoded in the models’ training data (Nadeem et al., 2021; Tan and Celis, 2019; Gehman et al., 2020).
In addition to generic capabilities, research has been carried out to investigate LLMs’ discipline-specific knowledge and skills. The inherent intricacies and nuances of specific domains necessitate tailored evaluation methods. For instance, researchers explored students’ and teachers’ usage of and perceptions toward LLMs, investigating opportunities and potential benefits for education (Kasneci et al., 2023; Baidoo-Anu and Owusu Ansah, 2023; Tlili et al., 2023). Researchers also tested LLMs’ ability to pass exams in specific domains like medicine (Gilson et al., 2023), law (Fei et al., 2023; Choi et al., 2023a), and business (Mollman, 2023). Others have evaluated LLMs’ ability to carry out real-world tasks. Examples include analyzing diagnostic abilities by answering questions about medications, symptoms and treatments (Lee et al., 2023); producing research hypotheses (Lahat et al., 2023); predicting legal judgment and recommending case citations (Chalkidis et al., 2019); and predicting chemical properties (Pan, 2023).
Many recent studies have explored the applicability of LLMs in the realm of data visualization. Inala et al. examine the potential impact of generative AI on the data analysis process, encompassing data collection, transformation, and exploration (Inala et al., 2024). DataFormulator 1 & 2 focus on facilitating data transformation workflows for creating data visualizations (Wang et al., 2023c, 2024c). On the other hand, several tools leverage LLMs to provide tailored natural language interfaces for generating data visualizations, including LIDA (Dibia, 2023), DynaVis (Vaithilingam et al., 2024), and ChartGPT (Tian et al., 2024). Other studies have evaluated LLMs’ capabilities in eliciting design preferences (Wang et al., 2024a), assisting in chart interpretation (Wang et al., 2024b), detecting misleading charts (Lo and Qu, 2024; Alexander et al., 2024), performing visualization literacy tasks (Bendeck and Stasko, 2024; Choe et al., 2024), generating visualization code (Chen et al., 2024), and troubleshooting visualization generation code (Shen et al., 2024). Furthermore, user studies have examined how ChatGPT is utilized in classrooms (Kim et al., 2024), visualization practitioners’ perceptions of text-to-image models (Schetinger et al., 2023), the impact of AI assistance on analysis workflows (Gu et al., 2024), and its effects on visualization coding performance (Vaithilingam et al., 2022).
The two pillars of data visualization skills are design and implementation. Most previous studies have focused on data visualization generation and implementation. In contrast, this work focuses on evaluating the design feedback capabilities of LLMs. Similar efforts have recently been explored in user interface design, where researchers have leveraged existing guidelines (Duan et al., 2024b) and critique datasets (Duan et al., 2024a) to perform design evaluations. In our study, we use real-world design questions from practitioners to conduct evaluations.
3. Assessing ChatGPT’s Competence in Data Visualization Knowledge
As a step toward assessing ChatGPT’s data visualization knowledge, we were interested in its ability to respond to real-world data visualization questions in comparison to Humans.
3.1. Data Collection
We used VisGuides (Diehl et al., 2018) as our main source for visualization questions. VisGuides is a discussion forum focused on visualization guidelines, where practitioners from various backgrounds, such as scientists, designers, engineers, and students, ask and respond to questions regarding their visualizations and fundamental visualization principles.
We identified a total of 226 questions within the VisGuides repository. Among these, we selected 119 questions for the study based on specific inclusion criteria. These criteria included the presence of a question in the post, the requirement for a visualization description to possess sufficient visual encoding (due to ChatGPT’s inability to process images at the time of our analysis), the need for questions to be comprehensible and well-structured (excluding overly generic inquiries like ’What do you think?’ or ’Is this good enough?’), and the necessity of having at least one Human response. Two of the authors jointly reviewed and unanimously approved the inclusion of these 119 questions.
We discovered that Visguides primarily featured two types of questions, broadly divided into two categories: design feedback questions (86) and visualization guideline questions (33). The former aimed to enhance users’ personal visualizations (see Figure 3 B & C), while the latter sought to comprehend visualization guidelines or principles (see Figure 3 A). For a more comprehensive taxonomy of questions on the platform, Diehl et al. provide detailed insights (Diehl et al., 2021).
After compiling the list of qualifying questions, we fed each query to ChatGPT 3.5 in May 2023 and with images to ChatGPT 4 with Vision in May 2024 (accompanied by images), accompanied by a role-playing prompt designed as follows:
Please act as if you are a visualization expert.
Can you respond to this question delimited by ///.
 {Question}
Please format your reply as JSON objects with keys: Question and Response.
    Question: xxx
    Response: xxx
There may be more than one question, format each as a new object.
To streamline our analysis, we created a spreadsheet where each row included a question, the corresponding ChatGPT responses (3.5 and 4V) and a compilation of Human responses. Each question had 1.54 Human responses on average ($Std.dev$ = 1.03). The average approximate word count of the Human responses was 235.35 ($Std.dev$ = 289.62) with a maximum of 1721 words and minimum of 13 words. The average approximate word count of the ChatGPT 3.5 responses was 463.42 ($Std.dev$ = 189.34) with a maximum of 1192 words and minimum of 115 words. For ChatGPT 4, the responses averaged around 255.03 words ($Std.dev$ = 88.58), with lengths spanning from 110 to 491 words.
3.2. Analysis Method
Two researchers performed an open coding process on a randomly selected 20% sample of the questions and the corresponding ChatGPT and Human responses. They iteratively agreed on and refined six evaluation metrics: coverage, topicality, breadth, depth, clarity, and actionability. Since design is a complex problem-solving task, we believe that qualitative human-centered metrics are more appropriate for capturing the holistic quality of responses compared to traditional utility metrics such as accuracy. These metrics were assessed using a Likert scale that ranged from 1: Very poor to 5: Very good. Table 1 presents our codebook of metric definitions and scoring rubrics.
Table 1. Metrics: Definitions and Scoring Rubrics
Metric & Definition Score & Description
Coverage assesses whether the response addresses all questions posed by the user. 1: Addresses 0–20% of the user’s questions.
2: Addresses 20–40% of the user’s questions.
3: Addresses 40–60% of the user’s questions.
4: Addresses 60–80% of the user’s questions.
5: Addresses 80–100% of the user’s questions.
Topicality indicates how closely the response stays on topic. 1: Mostly or entirely off-topic.
2: Partially addresses the topic but frequently digresses.
3: Somewhat relevant; minor deviations from the topic.
4: Mostly on-topic; occasional tangents.
5: Fully on-topic with no noticeable digressions.
Breadthmeasures how widely a response explores various ideas, concepts, options, or perspectives. 1: Very narrow; offering no alternative ideas.
2: Covers a small range of ideas, omitting key concepts.
3: Moderately diverse but limited in scope.
4: Addresses multiple angles with reasonable detail.
5: Very wide-ranging; covers a broad spectrum of ideas thoroughly.
Depth evaluates the extent of explanations, expertise, and valuable insights in a response. 1: Very minimal detail; lacks substance or justification.
2: Superficial treatment with limited insight.
3: Moderately detailed; some analysis or explanation.
4: Good explanation with solid reasoning and insight.
5: Deeper explanation, demonstrated with examples or expertise
Clarity assesses how easily a reader can understand the response, considering conciseness, organization, and readability. 1: Very unclear; difficult to follow or disorganized.
2: Somewhat unclear; main ideas partially obscured.
3: Moderately clear; understandable with some effort.
4: Clear, well-structured, and easy to read.
5: Flows seamlessly with minimal effort to comprehend.
Actionability assesses whether the response offers guidance readily implementable in a visualization (N/A for guideline-only questions). 1: No actionable guidance or next steps.
2: Little actionable guidance; mostly explanations.
3: Some actionable guidance; partially implementable.
4: Clear, actionable guidance that can be followed easily.
5: Highly actionable; provides explicit step-by-step instructions.
The researchers initially assigned scores to the 20% sample, engaging in discussions to resolve any discrepancies and establish consensus definitions and criteria for each score within every metric to ensure consistency. Following this, one of the researchers proceeded to score the remaining question responses. Throughout this coding process, they held continuous meetings to discuss any new issues and questions to ensure consistency and quality of codes and scores, refining metric definitions and score applicability.
Given that some questions elicited multiple Human responses while others received only one, we made them to be evaluated with a single score to assess the collective quality of the Human responses. This score captures the cumulative strengths of all sub-responses, recognizing that a single high-quality sub-response can adequately address the questioner’s needs even if others are less informative. We performed Mann–Whitney U tests to compare Human composite scores with ChatGPT response quality scores.
Furthermore, any elements of the Human responses that extended beyond the capabilities of ChatGPT, such as providing citations to related literature, links to external resources, visual examples, or posing follow-up questions, were recorded. We also categorized whether each question sought feedback on the questioner’s visualizations or inquired about general visualization design knowledge.
3.3. Analysis Decisions
Why Did We Use VisGuides?
We explored other sources of data visualization questions in the wild, including StackOverflow, the #help-general channel in the Slack workspace of The Data Visualization Society (dat, [n. d.]), and r/dataisbeautiful on Reddit. StackOverflow mainly focuses coding-related questions, which differ from our focus on design knowledge questions. The Data Visualization Society’s Slack channel, limited by a lack of a paid license, does not archive questions older than 90 days, and the queries are mostly technical, such as using specific software tools to create charts or visual effects. Reddit’s r/dataisbeautiful channel aligns more closely with our objectives, but a thorough review revealed that it contains more discussions and occasional critiques of web-found charts than actual questions. Hence, we decided to focus on the questions in the VisGuides forum.
Why Did We Choose ChatGPT Over Other Chatbots?
We selected ChatGPT as a representative of LLM-based chatbots for several reasons. While these chatbots use the same transformer architectures, ChatGPT is the most widely adopted in practice. Also, in our early pilot experimentation with Bard (now Gemini), Claude, and Bing (which uses the same GPT model as ChatGPT), ChatGPT generally provided responses that were better aligned in terms of style and content. For example, Bing often veered off by drawing information from various web sources, which were frequently inferior to ChatGPT’s internal knowledge responses. Furthermore, Bard’s recommendations were narrower in scope and often not applicable, like suggesting the addition of a trend line to a Gantt Chart. Overall, our impression was that these systems are mostly comparable, with ChatGPT having a slight edge over the others. Several studies also indicated that ChatGPT demonstrates superior performance compared to other LLMs in terms of a variety of task questions involving reasoning, facts, and fairness (Borji and Mohammadian, 2023; Ahmed et al., 2023). That said, we discuss an automatic evaluation pipeline, outlined in the discussion section, to enable comparisons of different models without requiring human intervention.
Why Did We Not Use Different Prompting Methods?
ChatGPT and other LLM-based chatbots are designed to follow human instructions (Wu et al., 2023; Liu et al., 2021b). Different prompting techniques, such as chain-of-thought (Wei et al., 2022b) and self-consistency (Wang et al., 2022), can lead to varied responses. To ensure a more objective evaluation, we used neutral prompts without any specific instructions (e.g., making the response concise) other than the role prompting. Moreover, we assume that users typically would not employ complex prompting strategies and would most likely use zero-shot queries. Thus, we merely forwarded the question from VisGuides to ChatGPT.
Why use The Six Quality Metrics?
Our goal was to assess the semantic quality of responses to visualization design questions—how well they conveyed meaningful, useful, and understandable knowledge—via human evaluation. While quantifiable measures (e.g., sentence structure, rhetorical style) can provide quick and repeatable approximations of text quality (Van der Lee et al., 2021), research has shown that they do not fully capture the complex and composite nature of constructs like diversity by quantifiable metrics (Van Miltenburg et al., 2018) and do not often correlate with how humans assess text quality (Scott and Moore, 2006; Reiter, 2018). Based on a review of literature on the human evaluation of automatically generated texts, we selected metrics that cover distinct aspects of intrinsic quality. Coverage and topicality capture the nuanced aspects of relevance—how the response answers the user’s questions directly (Van der Lee et al., 2021; Sreedhar et al., 2024; Awasthi et al., 2023); breadth and depth reflect the richness of ideas and reasoning (Wang et al., 2023b); actionability assesses practical applicability and informativeness (Van der Lee et al., 2021); and clarity is a widely used measure to assess readability, organization, and conciseness (Jiang et al., 2024; Awasthi et al., 2023). Overall, this framework offers a meaningful foundation for evaluating LLM performance in visualization design, with future work potentially extending it to include response tone, prioritization, or interaction quality.
3.4. Results
Figure 2. Mean scores (dots) and confidence intervals (error bars) across six metrics—Actionability, Breadth, Clarity, Coverage, Depth, and Topicality—for responses from ChatGPT 3.5, ChatGPT 4, and Human. Overall, ChatGPT 4 shows slightly higher mean scores than ChatGPT 3.5 and Human. Human responses exhibit higher variance but often perform comparably to ChatGPT.
Figure 3. Three selected and edited examples of questions and responses from the VisGuides forum. The ChatGPT responses consistently demonstrate breadth, topicality, and coverage. ChatGPT 4, in particular, showcases greater expertise by offering more in-depth and actionable advice in an authoritative manner, rather than merely listing ideas. The Human response ratings are more varied, as evident when comparing the first and last examples. Full questions and responses are provided in the supplement. Respondent IDs have been anonymized.
3.4.1. Ratings on Quality Metrics
Figure 2 shows the scores of response quality metrics for both ChatGPT and Human responses. We describe the findings for each quality metric as below.
Coverage
ChatGPT 4 achieved the highest score of 4.97 ($std.dev$ = 0.22), closely followed by ChatGPT 3.5 with 4.92 ($std.dev$ = 0.33). Humans scored lower at 4.37 ($std.dev$ = 0.95). Mann–Whitney U tests revealed that the difference between both ChatGPT models and Humans was statistically significant ($p<0.001$), while the comparison between ChatGPT 3.5 and ChatGPT 4 was not significant ($p=0.125$).
The lower score often stemmed from the Human respondent’s failure to address all facets of the question. For instance, in Figure 3C, the user raised a two-part question, “…would you recommend all the states be placed on the Y-axis instead of grouped as regions? If so, which visual layouts would be effective with dealing with the large number of states?…”. However, the Human response only addressed the second part of the question, suggesting a heat map as a useful visual layout.
Topicality
ChatGPT 4 achieved the highest score of 4.95 ($std.dev$ = 0.31), closely followed by ChatGPT 3.5 with 4.87 ($std.dev$ = 0.45). Humans scored slightly lower at 4.72 ($std.dev$ = 0.60). Mann–Whitney U tests revealed that ChatGPT 4 significantly outperformed Humans ($p<0.001$), and ChatGPT 3.5 also showed a statistically significant advantage over Humans ($p=0.014$). However, the comparison between ChatGPT 3.5 and ChatGPT 4 was not significant ($p=0.062$).
While all conditions performed well in addressing the correct question, there were moments when both ChatGPT and Humans strayed from the main topic. For instance, when a user inquired about the application of visual variables in VR visualizations,111visguides:visual-variables-in-vr ChatGPT 3.5 not only addressed this question but also came up with a hypothetical question asking what factors to consider when choosing such visual variables. Human respondents also occasionally veered off-topic by offering unsolicited advice and extending their responses beyond the original question’s scope.
Breadth
ChatGPT 3.5 achieved the highest score of 4.62 ($std.dev$ = 0.61), slightly outperforming ChatGPT 4, which scored 4.45 ($std.dev$ = 0.66). Humans scored significantly lower at 3.35 ($std.dev$ = 1.13). Mann–Whitney U tests revealed that both ChatGPT 3.5 and ChatGPT 4 significantly outperformed Humans ($p<0.001$ for both comparisons). Additionally, ChatGPT 3.5 demonstrated a statistically significant advantage over ChatGPT 4 ($p=0.025$).
For example, in Figure 3A, when the user inquired about color palette guidelines, ChatGPT provided an extensive list of potential approaches, encompassing data type, color harmony, accessibility, etc. In contrast, the Human respondent to this question concentrated primarily on the creation of domain-specific color maps, offering in-depth insights into the process.
Depth
For the depth metric, ChatGPT 4 achieved the highest score of 4.43 ($std.dev$ = 0.56), outperforming both ChatGPT 3.5, which scored 3.44 ($std.dev$ = 0.71), and Humans, who scored 3.52 ($std.dev$ = 1.07). Mann–Whitney U tests revealed that ChatGPT 4 significantly outperformed both ChatGPT 3.5 ($p<0.001$) and Humans ($p<0.001$). However, there was no significant difference between ChatGPT 3.5 and Humans ($p=0.518$).
There were intriguing and consistent behavioral patterns across the conditions. For example, ChatGPT 3.5 responses tended to provide general, surface-level improvements applicable to any visualization, such as adding labels and legends (Figure 3). While these suggestions were often hit or miss, they were effective when the questions sought generic advice. In contrast, ChatGPT 4 exhibited a more human-like behavior, offering focused suggestions in a more authoritative manner, showcasing expertise rather than merely enumerating a list of ideas, as illustrated in Figures 3 B and C.
Good Human responses were similar to ChatGPT 4, often even more nuanced, as demonstrated by the recommendation to “put a label before the numerical attribute like ‘Estimated Production: 1000’ (Figure 3B).” This detailed guidance directly pertains to the unique aspects of the visualization in question. However, Human responses also displayed greater variability. For instance, when asked about color clarity and alternative data representation to reduce clutter, a Human response simply confirmed the current design choices are fine222visguides:latitude-distribution-of-solar-and-wind-farms-uk. The lower scores of such poor quality single responses pulled the average down.
Clarity
Notably, ChatGPT consistently presented its responses in a well-structured list, which is a clear advantage of machine-generated responses. This contributed to ChatGPT 4 achieving the highest clarity score, averaging 4.97 ($std.dev$ = 0.20), closely followed by ChatGPT 3.5 with an average score of 4.93 ($std.dev$ = 0.28). Human responses also demonstrated good clarity, with an average score of 4.30 ($std.dev$ = 0.80). However, Humans received lower ratings when the responses were overly brief, making it difficult to discern specific reference points. Mann–Whitney U tests revealed that both ChatGPT models significantly outperformed Humans ($p<0.001$), while the difference between ChatGPT 4 and ChatGPT 3.5 was not statistically significant ($p=0.093$).
Actionability
ChatGPT 4 achieved the highest score, averaging 4.92 ($std.dev$ = 0.38), followed by ChatGPT 3.5 with an average score of 3.99 ($std.dev$ = 0.60). Human responses had the lowest average score of 3.85 ($std.dev$ = 0.94). Overall, ChatGPT 4 performed exceptionally well in providing actionable recommendations, as its responses consistently included focused, practical advice. By contrast, ChatGPT 3.5 often provided broader, less specific suggestions, which reduced its scores due to the generic or vague nature of the advice (Figures 3 B and C), although the list of recommendations often made sense (Figure 3A). The actionability scores for Human responses tend to be low when the users refrain from offering recommendations and instead validate all of the user’s choices333visguides:comparison-of-acute-hepatitis-from-tycho-dataset. Mann–Whitney U tests confirmed that ChatGPT 4 significantly outperformed both ChatGPT 3.5 ($p<0.001$) and Humans ($p<0.001$). However, there was no significant difference between ChatGPT 3.5 and Human responses ($p=0.416$).
3.4.2. Further Observations in Human and ChatGPT Responses
Humans tend to provide external resources
—43 out of 119 (36.13%). Human responses included elements that ChatGPT would not provide unless explicitly instructed; even when asked, it is known to frequently provide incorrect citations (Frosolini et al., 2023). This additional content encompassed references to academic research, links to related articles and websites, citations of studies, and the inclusion of informative video links. For instance, in Figure 3A, asking for guidelines on choosing color palette, the responses include references to three different academic research papers including “H. Fang, et al.”. In another case, when the question was about using blow-apart effects, the Human respondent embedded an educational video related to the subject444visguides:the-blow-apart-effect.
Multiple Human responses tend to complement each other
—Most questions had a single response (84), while 35 questions had multiple responses. Average ratings generally favored responses with multiple contributors across all categories, except for topicality. Significant rating differences were observed in coverage, breadth, depth, and actionability. Single-respondent questions averaged 4.25 ($std.dev$ = 1.02) in coverage, 3.06 ($std.dev$ = 1.05) in breadth, 3.35 ($std.dev$ = 1.05) in depth, and 3.72 ($std.dev$ = 0.96) in actionability. Conversely, multi-respondent questions scored higher, averaging 4.66 (*, $std.dev$ = 0.67) in coverage, 4.06 (**, $std.dev$ = 0.98) in breadth, 3.94 (**, $std.dev$ = 0.98) in depth, and 4.26 (*, $std.dev$ = 0.74) in actionability. When there’s only one response to a question, the quality of that response becomes heavily reliant on the individual respondent, leading to significant variability in quality, as seen in Figure 3A (high quality) versus Figure 3C (low quality). However, when multiple respondents contribute, they complement each other and compensate for areas in which one respondent might fall short (Figure 3B).
Human response rating improves for visualization guideline questions
—For ChatGPT responses, the differences in ratings between feedback and visualization guideline questions consistently remained small. For ChatGPT 4, there was a significant difference in the depth metric ($p=0.014$), with guideline questions receiving higher ratings (Mean = 4.64, $std.dev$ = 0.48) compared to feedback questions (Mean = 4.35, $std.dev$ = 0.57). For ChatGPT 3.5, the only significant difference was observed in the coverage metric ($p=0.002$), where feedback questions scored higher (Mean = 4.97, $std.dev$ = 0.24) compared to guideline questions (Mean = 4.79, $std.dev$ = 0.48). Across other metrics, no significant differences were detected, highlighting consistent performance regardless of the question type.
However, we observed more significant differences in Human responses between guideline and design feedback questions. For the breadth metric, guideline questions scored significantly higher (Mean = 4.12, $std.dev$ = 0.95) compared to feedback questions (Mean = 3.06, $std.dev$ = 1.05, $p<0.001$). Similarly, for the coverage metric, guideline questions achieved a higher score (Mean = 4.76, $std.dev$ = 0.43) than feedback questions (Mean = 4.22, $std.dev$ = 1.05, $p=0.015$). The depth metric also showed a notable difference, with guideline questions receiving a substantially higher score (Mean = 4.24, $std.dev$ = 0.95) compared to feedback questions (Mean = 3.24, $std.dev$ = 0.98, $p<0.001$). Overall, Human responses to guideline questions tended to be more comprehensive and detailed. For other metrics, no significant differences were observed.
Guideline questions frequently require broader perspectives and readily accessible design knowledge, with 84.4% (27/32) of Human responses citing external references, in contrast to design feedback questions, where only 18.39% (16/87) of Human responses did so. Conversely, feedback questions entail an understanding of domain-specific data and tasks, potentially making it challenging to offer comprehensive insights. For instance, a question like “Is this color scheme suitable?” requires a deep understanding of the domain. Conversely, questions such as the one about gendered colors in visualizations555visguides:use-of-gendered-colours-in-visualization can easily elicit a variety of various viewpoints and existing resources.
Question specificity contributes to response variability
—While we did not quantify this, we observed ChatGPT’s response might be susceptible to how specific and clear the question was. Figure 3B shows such an example when the user query was “do you believe the graph is clear?” Other examples of less specific user questions include: “is my color map optimal?” and “how can my visual design be improved?”666visguides:map-and-bar-visualization. Such commonly encountered ill-specified questions can pose a challenge for LLMs, as they may begin suggesting irrelevant feedback instead of asking clarifying questions. On the other hand, in Figure 3C, the question is more specific, resulting in more meaningful options and insights offered by ChatGPT.
ChatGPT 3.5’s color vision deficiency was usually not problematic—Color-related questions were prominent. Out of the 119 questions, 32 were centered on color-based design feedback (e.g., “Is the choice of the colour scheme appropriate?” 777visguides:area-chart-generation-capacity-over-the-years). Users rarely mentioned the specific colors used in their visualizations, instead providing only visual encoding information in their questions. Consequently, ChatGPT 3.5 lacked access to exact color information due to its inability to comprehend images. Despite this limitation, there were no significant differences across all metrics in ChatGPT 3.5’s responses.
3.5. Takeaways
Our analysis shed light on ChatGPT’s cability to respond to data visualization queries compared to Human counterparts. ChatGPT clearly excelled in breadth, exploring a wide range of ideas and concepts, surpassing Human responses. Clarity was another area where ChatGPT scored highly, benefiting from its structured response format. ChatGPT 4 outperformed ChatGPT 3.5 by providing more focused and in-depth responses, aided by its image comprehension capabilities. Some Human responses resembled those of ChatGPT 4, often enhancing their quality by including reference resources and complementing one another in multi-respondent scenarios.
Participant ID Role Years in Data Vis Gender Age range Racial background Frequency of ChatGPT usage
P10 Developer 3-5 years Male 25-34 years old Asian Daily
P11 Manager 1-3 years Male 25-34 years old Caucasian Occasionally
P12 Freelancer 3-5 years Male 35-44 years old Asian Daily
P13 Journalist 1-3 years Female 18-24 years old Hispanic Weekly
P14 Consultant 3-5 years Male 25-34 years old Caucasian Daily
P15 Product Designer 3-5 years Female 25-34 years old Caucasian Weekly
P16 Analyst 5-10 years Male 35-44 years old Caucasian, African American Weekly
P17 Scientist 3-5 years Female 25-34 years old Asian Weekly
P18 Scientist 1-3 years Female 25-34 years old African American Weekly
P19 Student 5-10 years Male 25-34 years old Caucasian Only once or twice
P20 Freelancer ¿ 10 years Male 45-54 years old Caucasian Occasionally
P21 Student 3-5 years Male 25-34 years old Asian Daily
Table 2. Demographic and experience related information for participants in user study.
4. Perception of practitioners toward ChatGPT’s utility
While the VisGuides forum analysis findings are promising, they do not provide insights into how practitioners perceive the value of ChatGPT’s responses, prompting the need for further research. As a result, we conducted a comparative interview study involving feedback sessions with ChatGPT and Human experts. Experts were chosen over anonymous users to ensure the quality of human responses and establish a higher benchmark for evaluating the limitations of ChatGPT. These experts also represent experienced colleagues who provide design feedback in professional settings (Choi et al., 2023b; Esteves and Neves, 2022). Our goal was to explore how data visualization practitioners evaluate ChatGPT’s effectiveness as a design companion, focusing on their experiences and perspectives on its design feedback.
4.1. Recruitment
We aimed to engage a diverse group of data visualization practitioners in our study. We recruited participants through academic mailing lists targeting students and scientists, as well as the Data Visualization Society’s Slack channel (dat, [n. d.]). Inclusion criteria required proficiency in English, experience in data visualization creation, and a willingness to share at least one of their data visualizations. Initially, 41 individuals responded to the recruitment survey, from which we selected 12 participants to take part in the study.
4.2. Participants
The study involved 12 participants (P10 to P21), creating a diverse group with varying professional backgrounds, experience levels, and familiarity with ChatGPT (see Table 2); we started participant codes from 10 to ensure that all IDs were consistently two digits long. Their roles encompassed a broad spectrum, including developers, managers, journalists, consultants, product designers, analysts, scientists, and students. Professional experience ranged from less than one year to over ten years in their respective fields. Furthermore, participants’ engagement with ChatGPT varied, with some being daily users while others interacted with it only occasionally or as the need arose.
4.3. Tasks & Procedures
Before each study session, participants were asked to prepare a visualization they were comfortable sharing, along with a list of relevant questions. These sessions were conducted using Zoom between the end of July and early August 2023. The entire study session took about 60 minutes. Participants were compensated with a \$50 Amazon gift card.
Each interview began with a brief introduction by the moderator and was divided into three segments: a visualization feedback session with ChatGPT, a similar session with Human expert(s), and an open-ended interview. Our study employed a within-subjects design with counterbalancing to control for order effects. Six participants initiated the process with the ChatGPT session, while the other six commenced with the Human expert session. After each feedback session, we administered a survey. In five sessions, two visualization experts provided feedback, while in the remaining seven sessions, we had one expert present. Both experts are current professors in the field of visualization.
During the ChatGPT feedback session, participants shared their screens with the moderator and presented their visualizations, accompanied by a brief explanation. The moderator introduced a predefined input format for ChatGPT, employing a role-playing structure (Ihwan, [n. d.]). Participants were guided to furnish details about the visualization’s chart type, textual description, visual encodings, and any related questions. They input this information into ChatGPT while screen-sharing. Following ChatGPT’s response, participants could pose follow-up questions or request clarifications. Afterward, participants received a survey via the Zoom chat to gather feedback on their ChatGPT experience.
During the Human expert session, one or two visualization experts joined the Zoom meeting. Participants shared their screens and presented their visualization-related questions to these experts. The experts answered queries and provided additional feedback or further insights. Subsequently, following this session, the visualization experts left the meeting, and participants were presented with a Zoom survey to gather feedback about their experience.
In the final segment of the study, open-ended interviews were conducted. In these interviews, the moderator inquired about participants’ experiences in both ChatGPT and Human feedback sessions, delving deeper into the rationale behind their survey responses. Furthermore, participants were prompted to share their perspectives on the possible integration of ChatGPT into their data visualization workflow, its constraints, and the future possibilities it might offer.
4.4. Data & Analysis Methods
The final dataset from our user studies comprises participant questions (see Table 3) and responses, survey data, and transcripts from post interviews. Our primary focus for analysis was on the interview transcripts. Initially, two researchers independently examined three interview transcripts, identifying interesting quotes and assigning meaningful themes. Once we established a high degree of consistency between the researchers’ findings, one of them proceeded to review the remaining transcripts. Ultimately, we compiled a categorized list of quotes based on initial higher-level themes (e.g., Where Human experts excel). To further refine our analysis, we revisited these quotes and divided the themes into sub-level categories (e.g., Collaborative and natural conversations).
4.5. Post-Session Interview Analysis Results
P ID Chart Type Initial Questions
P10 parallel coordinate map Could you suggest an alternative way to visualize this? How can I make this visualization more engaging?
P11 scatterplot Does it make more sense to compare X or to compare Y? Would there be a better visualization type to show ___?
P12 scatterplot How can I improve this visualization in order to satisfy the customer?
P13 scatterplot What is the best way to show ___? what do you think about the current color scheme?
P14 diverging stacked bar chart How can we ensure the user looks at ___? How can I put these four bars next to each other in one visualization?
P15 bubble cluster How would I enable more properties to be seen? How would I let the user encode the strength of their preference?
P16 interactive map Which parts of the data visualization do you think are successful? Are there any areas you believe could be improved for better clarity and impact?
P17 sankey chart Is there a way to ascertain what visualization is most effective for communicating the data? How could ___ be more clearly displayed?
P18 wav file visualization Would it be visually overwhelming to use ___? Is this an appropriate visualization for demonstrating ___?
P19 color heatmaps Would another color map be more suitable? Would you have any suggestions as to how to indicate ___?
P20 beeswarm, line chart Does the special encoding for ___ make intuitive sense? Is it clear that ___ represents ___?
P21 comparative line graph How can I decrease the number of colors used in this visualization? Would this data be better represented if I ___?
Table 3. Initial questions from participants. These initial questions are altered to maintain the privacy of the participants. The original questions provided more contextual information and specificity.
4.5.1. Participant Questions and Interaction Dynamics
Four of the participants (P10, P15, P16, and P18) shared interactive visualizations. The other eight participants shared static visualizations. The type of visualization, as well as the questions they initially brought, can be seen in Table 3. To safeguard participant privacy, we refrain from disclosing their visualizations or any elements of their questions that could potentially reveal information about participants.
In the ChatGPT section of the interview, the majority of the participants asked ChatGPT more than one follow-up question. The follow-up questions to ChatGPT tended to be more specific and technical than the follow-up questions asked to the Human experts. For example, P11 asked ChatGPT “What are the pros and cons of having too many bubbles on the scatterplot from having a clear comparison?”
On the other hand, the follow up questions to the Human experts tended to be less structured and more conversational. The participant and expert had a conversation and pulled out different aspects of the question along the way. There was a lot of stopping to clarify what different elements of the visualization meant and how they played into the visualization. As a result, other factors not explicitly mentioned in the question often ended up getting brought up in these sessions with the Human experts.
4.5.2. Post Survey Results
In Figure 4, we present the outcomes of the experience surveys conducted following each feedback session. The overarching consensus among participants was a strong preference for Human experts. This preference was underscored by statistically significant distinctions, as determined through two-tailed Mann-Whitney U tests (1: strongly disagree to 5: strongly agree). Specifically, participants expressed significantly higher levels of satisfaction in interacting with Human experts ($U=113.5$, $p<0.05$), perceiving their responses as notably more accurate ($U=113.5$, $p<0.05$), helpful ($U=110.0$, $p<0.05$), reliable ($U=128.5$, $p<0.05$), and adaptable ($U=105.5$, $p<0.05$) to their preferences.
Furthermore, respondents indicated that Human experts exhibited a significantly deeper understanding of the context and requirements of their queries. Human experts were also acknowledged for their expertise in data visualization, along with their ability to offer actionable recommendations. It is worth noting that no significant differences were observed in responses to questions related to the clarity and conciseness of explanations, concerns about potential biases, or perceived risks of receiving misleading information.
Figure 4. User preference for Human expert vs. ChatGPT responses in feedback sessions: The overarching pattern evident in these charts predominantly indicates a user preference for Human experts over ChatGPT in the context of feedback sessions.
4.5.3. Where Human Experts Excel
Participants were generally more satisfied with Human experts. They mentioned a variety of reasons for such opinions. The participants highlighted several key themes in their responses, shedding light on the strengths of Human experts in the context of data visualization guidance.
Bespoke and focused feedback of Human experts
First of all, they liked Human experts’ ability to offer tailored recommendations that were closely aligned with the specific visuals in question (P10, P11, P15, P17 - P21). P17 exemplified this sentiment by stating, “[…] recommendations are a lot more tailored and specific in that way, versus just like being tossed a list of potential tools we can look.”. P11 further elaborated on this point by highlighting that Human experts can envision various options and present the ones they deem most valuable. They also noted that Human experts could effectively understand the context of visuals by directly observing them (P11, P17 - P19). P19 expressed this sentiment by stating, “I felt the feedback was obviously more grounded to the visualization at hand. Because again, they could see it.” Similarly, P18 commented, “As he’s looking at the pictures, he knows the information I had access to and how that translates.” Furthermore, participants acknowledged that Human experts excelled in staying focused on the specific problem at hand. For instance, P15 articulated, “[the expert] was able to talk about improvements within scope […] without redesigning it entirely.” Similarly, P10 commented, “The ones that [the expert] gave were within the scope of refining existing visualizations.”
Collaborative and natural conversations
Participants consistently expressed the value of more fluid conversations when interacting with Human experts (P10, P11, P15, P17, P19, P20). They highlighted that Human responses felt more free-flowing, contributing to a sense of genuine interaction (P10, P17). Participants also noted that Human-expert interactions provided a collaborative and interactive experience (P11, P15, P19, P20). For instance, P11 highlighted an efficient turn-taking dynamic by saying, “the ability to ask follow-up questions and clarify I think was easier to do within the Human interaction context.” Others similarly emphasized the convenience of seeking clarifications without the need for precise phrasing or additional context adjustments (P11, P15).
Enriched insights through lived experience
Participants felt that Human experts’ responses were often rooted in their experience and education (P16). They valued the ability of Human experts to explain the underlying rationale, promoting a more actionable approach to implementing the suggestions (P10). Furthermore, participants appreciated the Human experts’ ability to think beyond the obvious; as P13 noted, “I feel like I was definitely more satisfied with the Human experts, […] they just started thinking so many things beyond anything I could have think of.” Moreover, participants conveyed that interactions with Human experts felt personal and supportive. One participant reflected, “they were really walking me through the different things […] help you grow as a professional […] really pushing me.”, while another participant recognized the proactiveness of Human experts by saying “even if I didn’t ask question, they kind of brought it up.”
4.5.4. Where ChatGPT excels
Participants identified the strengths of ChatGPT in contrast to Human experts.
Brainstorming and ideation
ChatGPT’s role as a creative catalyst emerged prominently during participant discussions, underlining its capacity to spark innovative ideas (P11, P15, P18). Participants like P18 found unexpected and impressive suggestions, stating, “It did have some cool ideas, I didn’t expect this from ChatGPT. […] Pretty good.”, while P11 commented, “I think it’s really interesting to try and kind of break outside the box and get more creative… GPT has the potential to disrupt that workflow in a positive way.” Likewise, P15 also acknowledged that the tool may not always provide precise solutions but rather serve as a springboard for creative exploration. P15 contrasts this to Human experts by saying, “[expert] feedback was less brainstorming and more actual critique.”
Broad knowledge base
In a similar vein to its brainstorming capability, participants also highlighted ChatGPT’s capacity to serve as a vast repository of knowledge, offering a wide array of ideas (P10, P11, P15). P10 stated, “it really does kind of feel like a really fancy kind of search engine, where it is kinda, I give it a problem. And it’s like, oh, like, this might be a solution.”, while P15 said, “great place to go to start […] researching or come up with ideas you hadn’t thought of.” P11 similarly emphasized that ChatGPT excels at showcasing a range of data visualization options, making it a valuable resource during the initial stages of projects. P11 elaborated on the tool’s potential, envisioning a scenario where metadata about a dataset could be fed to ChatGPT to receive chart-type recommendations. However, P10 emphasized that a critical eye is needed to discern the relevant and valuable insights from its outputs; “I asked to give 20 ideas and I’m […] experienced enough […] to […] know that, like, 16 of them are just nonsense. The other four things I might not have thought of.”
Time saved via rapid understanding and response
Participants appreciated ChatGPT’s ability to quickly grasp and provide information (P10, P11, P13). P10 said ChatGPT was able to quickly grasp unfamiliar concepts auxiliary to visualization (e.g., explaining Wordle to ChatGPT), while P11 and P13 felt ChatGPT provided faster and more comprehensive lists of options, especially for alternative chart types, when compared to the responses from Human experts. On a related note, P11 said that Human interactions can sometimes involve a lengthier back-and-forth process to arrive at the right question.
Soft attitude and neutral perspective
Other participants noted ChatGPT’s behavioral traits. For example, P19 pointed out that ChatGPT often assumed a gentle and agreeable attitude by acknowledging concerns without offering specific directions, which led to a perception of excessive alignment. P19 also noted that ChatGPT demonstrated a neutral stance, presenting a variety of viewpoints without favoring any particular side. This contrasted with Human experts who might convey their own biases, thereby making ChatGPT’s approach seem more balanced.
4.5.5. Where ChatGPT falls short
Participants also discussed ChatGPT’s several shortcomings in comparison to Human experts.
Lack of depth in responses
Participant feedback indicated a notable lack of depth in ChatGPT’s responses (P13, P15, P19). For instance, P19 expressed that ChatGPT’s advice seemed to lack actionable insights and depth that Human experts possess. This sentiment was echoed by P15, who noted that while ChatGPT’s answers were broad in knowledge, they lacked depth. P15 further elaborated that ChatGPT often suggested radical changes (i.e., diverse ideas), while Human experts provided improvements that were more aligned with refining existing visualizations. P19 similarly commented that ChatGPT’s ideas were not considering the specific visualization. Participants acknowledged this contrast as a trade-off (P15, P17), emphasizing that the desired level of response detail and granularity from either ChatGPT or Human experts depends on user needs.
Generic advice, misalignment, and lack of critical thinking
Participants expressed that ChatGPT’s feedback often felt generic and lacked contextual depth (P12, P16, P20). P20 expressed that ChatGPT’s “knowledge felt […] not necessarily backed up by experience.” P16 and P12 similarly noted that ChatGPT’s responses were not necessarily wrong but were overly generic, failing to consider the actual outcomes of the provided recommendations. P16 further commented that they “felt the areas of improvement were good suggestions. But it didn’t feel like you have to do this.”
Several participants highlighted instances of misalignment and a lack of specificity in ChatGPT’s recommendations (P11, P16, P19). P19 pointed out an example where ChatGPT suggested color maps that might create confusion rather than improve visualizations. This lack of alignment was further emphasized by P11, who noted a communication hurdle hindering the establishment of a common understanding and precise feedback.
Participants highlighted concerns related to critical thinking (P13, P16). P16 noted, “whereas the ChatGPT felt more like it was telling me what I wanted to hear if that makes sense.” P16 further elaborated that ChatGPT’s responses sometimes resembled textbook information rather than offering insightful suggestions on visualization. P13 shared a similar sentiment, stating, “And I feel like the issue with ChatGPT typically is that it only responds to your question. So if you cannot think of the question, it will not give you the answer.”
Efforts required for fluid conversation
Participants highlighted the challenges in having a fluid and interactive conversation with ChatGPT (P10, P13, P16, P17, P20). For instance, P10 explained, “… kind of hoping that one of them would be what I wanted.”. P17 expressed a similar sentiment, stating, “even though it’s still conversational, but it still kind of feels a little bit static sometimes where you’re just typing text into a text box.” P16 further emphasized the difference in interaction dynamics, stating, “I really liked listening to the Human experts go back and forth, and talk about different aspects… whereas ChatGPT just fed that to me, and that was it.” P13 noted that ChatGPT’s responses often required iterations and follow-ups, P10 found that ChatGPT’s insights became more specific and interesting when more constraints were provided.
4.5.6. Opinions on Trust and Reliability with ChatGPT
Participants had varied opinions on the trustworthiness and reliability of ChatGPT’s responses, with some expressing low trust (P10, 18, P20). P20 mentioned the limitations in ChatGPT’s understanding of visuals, which relied heavily on the accuracy of participants’ descriptions. P18 pointed out the potential for fabricated information due to the algorithmic nature of the responses. P10 echoed this sentiment, indicating that ChatGPT often sounded too confident and authoritative, which could mislead users into placing unwarranted trust. On the other hand, P17 said the two are relatively the same, highlighting that the key distinction lies in how information is delivered.
Several other participants pointed to the need for due diligence in evaluating ChatGPT’s suggestions, especially when the user lacks domain expertise (P10, P15). For instance, P10 commented: “If I was in a field that I wasn’t familiar with, I think it’d be really easy to get fooled by it.” P15 drew a parallel with the early inception of Wikipedia, stating, “ it was like early days of Wikipedia, I was taught like, never cite Wikipedia, it could be wrong. And I think the same thing is true of something like a large language model.”
They also made comparisons to Human experts in terms of trust and authority (P13, P16, P18, P21). They explained how the credentials of Human respondents influenced their perception (P18, P21), e.g., “Well, the Human respondent is a professor in visualization.”—P18. Others perceived Human experts as sounding knowledgeable and educational (P13, P16). In contrast to these, participants also shared their perceptions of expertise and self-awareness in AI. P16 said, “Didn’t feel like that the ChatGPT response was incorrect … trusted it, but not nearly as much as the Humans.”, while P20 expressed, ChatGPT doesn’t know when it’s wrong.”
4.5.7. Limitations and Opportunities with ChatGPT
Participants expressed an optimistic outlook regarding the future potential of ChatGPT, despite acknowledging its current challenges (P10, P12, P15, P16). They discussed aspects they would like to see improved in ChatGPT in the future.
Ability to convey complex visualizations
Participants emphasized the significance of enabling ChatGPT to comprehend and interpret complex visualizations, highlighting the limitations of text-based communication when dealing with intricate design problems (P10, P11, P13, P15 - 21). P15 acknowledged the inherent loss of information when translating between visual and written mediums, while P10 and P11 conveyed the complexity of describing interactive dashboards and multimedia visualizations via text. P10 also mentioned video recordings that could convey the interactivity of designs.
Ability to generate visualizations based on feedback
Several participants pondered the prospect of ChatGPT being able to not only analyze input information but also generate relevant visualizations (P18, P19, P21). For example, P18 proposed the generation or modification of images based on textual descriptions. P19 noted the convenience of sharing underlying code to accompany visual content, emphasizing its value in effectively communicating complex ideas instead of relying solely on text descriptions. Similarly, P21 discussed the concept of inputting data and posing questions to explore the space of chart design. P13 envisioned the integration of ChatGPT within existing tools, enhancing workflow efficiency and facilitating a more cohesive user experience.
Facilitating Fluid and Truthful Conversations
Participants expressed the need for ChatGPT to initiate and support more fluid conversations to enhance its usability and value (P11, P16, P21). P16 highlighted the need for AI to prompt users for further details about the visualization; “I would really like it if it were to ask me questions to kind of gauge my thought process.” Similarly, P11 noted the value of follow-up questions in driving valuable insights, while “it was harder to envision what a follow-up question to that system would be.” On the other hand, P21 raised concerns about the possibility of ChatGPT providing false or hallucinated information.
The indispensable role of Human feedback
The consensus among participants highlighted the enduring value of Human feedback in evaluating and improving visualizations (P11, P16, P21). P16 commented, “I think there’s always, at least in my lifetime, going to be a need for Human interpretation. And the Human experts’ sessions, for me, reinforce that they offer a lot of really good insight.” P11 further expressed that despite ChatGPT’s capabilities, the feedback from actual users remains incomparable, as it encompasses a deeper level of insight into user experience. Additionally, P11 discussed ideas of experimenting with different personas for hypothetical user interactions.
Issues with privacy and sensitive data
Participants in the interviews expressed significant concerns regarding the privacy and handling of sensitive data by ChatGPT (P15, P16, P20, P21). They emphasized that sharing visuals containing private or sensitive information would be uncomfortable and unlikely unless the data was from a public source or there was a guarantee that it would not be shared publicly. P15 shared an illustrative example, stating, “I met a company that I had to sign an NDA with, and we are not allowed to use ChatGPT with anything that might be considered confidential information because that information goes onto OpenAI servers and will live there for eternity.”
4.6. Takeaways
The findings from our study underscore the distinct strengths and limitations of ChatGPT in comparison to Human experts. The clear preference for Human experts among participants highlights their superior ability to provide tailored, context-specific advice and engage in more interactive and collaborative conversations. This preference is rooted in the Human experts’ deep understanding of visualization context and their capacity for nuanced, bespoke feedback. The feedback provided by Human experts in these sessions closely paralleled the high-scoring Human responses observed in the Visguides.
On the other hand, ChatGPT shows promise as a creative catalyst and a broad knowledge base, offering rapid responses and a diverse array of ideas. However, its limitations are evident in the lack of depth and specificity in responses, challenges in fluid conversation, and concerns about perceived trustworthiness (Yin et al., 2019; Toreini et al., 2020; Drozdal et al., 2020; Yang et al., 2020; Chatzimparmpas et al., 2020) and handling sensitive data (Liu et al., 2021a). Several participants expressed concerns about needing expertise to assess the validity and relevance of potentially inaccurate suggestions from ChatGPT. Nonetheless, they remained optimistic about its future potential, anticipating continued improvements and increasing value.
5. Discussions
5.1. Limitations of our studies and future opportunities
While our studies were designed and conducted with considerable care, there exist certain limitations. First, the absence of demographic and expertise data for the Human respondents in the VisGuides forum makes direct comparisons with the Human experts challenging. However, the human baselines in both studies represent distinct real-world scenarios: receiving feedback from anonymous online platforms or from experienced colleagues in the workplace. Also, due to the dynamic nature of the feedback sessions in the second study, we were not able to apply the static evaluation metrics used in the VisGuides. Despite this limitation, our findings were highly consistent across the two studies.
We used a fixed version of ChatGPT. ChatGPT and other LLM-based chatbots are evolving to incorporate more recent data and support new features, including the ability to handle more nuanced language or to understand and generate images, although some of these features are only available to paid users. As a result, responses to the questions in our study can vary among newer LLMs. Likewise, the recently released advanced voice mode could help narrow the interaction modality gap identified in the second study, potentially improving the fluidity of conversations. That said, it still falls short of achieving human-like naturalness, such as exhibiting proactive questioning and avoiding passive behavior. However, our study results provide a valuable foundation for future research, offering both datasets, metrics, and methodologies for evaluating LLM responses in the context of visualization design feedback.
5.2. Towards automatic evaluation of LLM feedback capabilities
Facilitating ongoing monitoring of LLMs’ design feedback capabilities would be beneficial to ensure that both the development of LLMs and the systems built upon them remain well-grounded. The current reliance on human intervention to evaluate the quality of responses is inefficient. Recently, LLMs have demonstrated remarkable capabilities in evaluating natural language generation, showing greater alignment with human evaluations compared to traditional similarity-based or embedding-based metrics such as BLEU or BERTScore (Zheng et al., 2023; Liu et al., 2023b; Wang et al., 2023a; Fu et al., 2023; Kim et al., 2023).
Metric Response type Mean diff $\tau$ $\rho$ Gwet’s AC1 Score dist
Depth Human -0.563 0.558 0.631 0.208
ChatGPT 0.542 0.215 0.228 0.29
Breadth Human -0.24 0.553 0.62 0.424
ChatGPT -0.083 0.522 0.537 0.713
Clarity Human -0.792 0.372 0.395 0.057
ChatGPT -0.01 0.517 0.52 0.901
Topicality Human -0.771 0.366 0.396 0.208
ChatGPT 0.063 -0.029 -0.029 0.915
Actionability Human -0.928 0.47 0.525 0.167
ChatGPT 0.406 0.466 0.492 0.422
Coverage Human -0.87 0.41 0.457 0.097
ChatGPT 0.01 0.424 0.424 0.925
All All -0.27 0.404 0.433 0.444
Human -0.694 0.455 0.504 0.194
ChatGPT 0.154 0.353 0.362 0.694
Table 4. The experiment results of the LLM-based automatic evaluation. The alignment between scores from LLM and human evaluators was derived from mean difference (Human score - LLM score), and the Kendall’s tau ($\tau$), Spearman ($\rho$), and Gwet’s AC1 (AC1) correlation coefficients. The histograms show the score distribution ranging from 1 (left) to 5 (right) for each metric.
We tested the viability of this idea by developing an evaluation prompt (see below) and used it to assess both Human and ChatGPT 3.5’s responses based on the six evaluation metrics outlined in Table 1.
In the experiment, we employed the in-context learning method (Brown et al., 2020; Li et al., 2024) using GPT-4o-mini via API calls in December 2024, leveraging the LLM’s few-shot learning capabilities to generalize evaluation tasks from the provided examples. Based on these settings, the evaluation task was conducted in the following steps: First, the Visguides dataset was split into a few-shot development set (20%) and a test set (80%). Second, for each test case and a metric, $n$ examples were selected from the few-shot development set, where we settled on $n=5$ after testing values within a range between 1 to 8 given the limit of context lengths in the GPT-4o-mini. The process for the example selection was based on two principles: 1) Score pooling for extracting diverse examples: To ensure a diverse pool of examples ranging from low to high quality, the examples in the few-shot development set were grouped into bins for unique scores from 1 to 5 based on their scores for the target metric. 2) Embedding similarity for retrieving relevant examples: Within each bin, the most similar examples based on embedding similarity based on Sentence Transformers (Reimers and Gurevych, 2019) were selected to help guide the model to choose the examples that are most semantically close to the input query, effectively “picking the best match” from the limited example set. The number of examples was determined proportionally to the size of the corresponding bin. The selected examples were incorporated into the prompt and provided to GPT-4o-mini, released in August 2024, as presented below:
You are an expert evaluator specializing in data visualization assessments.
Your task is to evaluate responses to visualization-related questions as closely
aligned with given examples, focusing specifically on the {metric} metric.
Here are specific guidelines for {metric}:
    Before evaluating, carefully consider:
    1. The specific visualization requirements in the question
    2. How well the response addresses those requirements
    3. The technical accuracy and appropriateness of the visualization approach
    4. The practicality and implementability of the solution
   {metric-specific evaluation rubric}
   {metric-specific evaluation guideline}
Here are some example evaluations:
   {few-shot examples}
Now evaluate this {response type} response based on the following metric:
    Question: {question}
    Response to evaluate: {response}
Evaluate the response on {metric} using a scale of 1-5 (1 being lowest, 5 being
highest). Your job is to make your evaluation as closely aligned as the given examples,
calibrating the score closely to the examples.
Metric definition:
- {metric}: {metric definition}
For the criterion, provide:
1. A numerical score (1-5)
2. A brief justification for the score
3. Specific examples from the response
Your response MUST be a valid JSON object with this exact structure:
    "evaluation": {
        "{metric}": {
            "score": X,
            "justification": "...",
            "examples": "..."
In the prompt, the metric-specific information such as evaluation rubric and guideline (Table LABEL:tab:eval-rubrics and LABEL:tab:eval-guidelines), few-shot examples, metric definition (Table 1), and response type (Human or ChatGPT) are integrated for each test case accordingly.
We then compare the output scores to those provided by human evaluators and calculate the alignment using agreement measures and correlation coefficients. In our experiment, we use Kendall’s tau ($\tau$), Spearman ($\rho$), and Gwet’s AC1 coefficient as underlying measures. While Kendall’s tau was selected as one of the most widely used metrics, several studies (Hou, 2024; Liu et al., 2023a) have pointed out that this measure can lead to biased results when the distribution of the underlying scores is severely skewed, failing to capture the true alignment. This issue is particularly relevant in our case for metrics such as clarity, coverage, and topicality, especially for ChatGPT’s responses, where scores are skewed toward higher values (e.g., 4 or 5). To address this, Gwet’s AC1 (Gwet, 2001), which has been proposed and increasingly adopted to mitigate biased judgments, was used to complement Kendall’s tau.
Table 4 presents the outcomes of these measures for each evaluation metric. Overall, the results show that the LLM evaluator’s scores are fairly aligned with those of humans ($\tau=0.404$, $\rho=0.433$) obtained across all metrics, given a well-known interpretation of Kendall’s tau and Spearman coefficient (Akoglu, 2018; Dancey, 2007) ($\pm$0.4-0.6: moderate agreement). Specifically, the alignment was better captured for both Human and ChatGPT responses in the breadth ($\tau=0.553,\rho=0.62$; $\tau=0.522,\rho=0.637$) and actionability metric ($\tau=0.47,\rho=0.525$; $\tau=0.466,\rho=0.492$), and in the depth metric for Human responses ($\tau=0.558,\rho=0.631$). For metrics with highly skewed score distributions, such as topicality, Kendall’s tau coefficients were relatively low for ChatGPT responses. In these cases, Gwet’s AC1 coefficient provided a more nuanced perspective, demonstrating a high alignment ($\tau=-0.029,\rho=-0.029,AC1=0.915$). While our experiment shows promising capabilities of automatic evaluation using LLMs, it is worth to further investigate refining heuristics and developing prompts, for example, making prompts targeted to response types or providing clearer metric descriptions, in order to improve the alignment.
6. Design considerations for LLM-based feedback systems
Our study results indicate the feasibility of developing an advanced design feedback system using LLMs. Drawing from the insights gained, we address several key design considerations for constructing such a system by addressing the strengths and limitations we identified in the study.
Leverage the knowledge base for creative exploration
It was evident that the broad knowledge base of LLMs is advantageous for rapidly exploring diverse ideas or sparking new ideas for participants. However, one issue we observed in the study was that this discovery process was mostly ad-hoc and spontaneous, resulting in a hit-or-miss experience. Future feedback systems have the potential to make this process more systematic and supportive of structured ideation. This could be achieved by providing prompts or guiding questions that steer exploration toward relevant and productive paths. For example, a recent study provided insights into a structured exploration of the design space by generating dimensions to guide response generation (Suh et al., 2023).
Support nuanced understanding of design context
One of the most frequent complaints was the lack of a nuanced understanding of the problem context in the question. LLMs with enhanced image understanding capability will alleviate this issue but will not be enough. Future systems might consider allowing visual prompting in which users can annotate areas of interest. Participants also mentioned generating image responses that can help users envision how to address their design concerns; i.e., these responses could also be annotated to highlight improvements. Moreover, communicating design issues about interactive data visualizations poses interesting research challenges for future systems.
Facilitate fluid and proactive engagements
Our participants favored seamless and dynamic discussions with Human experts. ChatGPT is currently trained to align with users’ intentions (Ouyang et al., 2022), often resulting in passive and affirmative behavior. Future feedback systems can aim to be more proactive, especially for novice users who may struggle to articulate their design concerns, such as by asking clarification questions or suggesting related design questions. Participant comments indicate that this approach may involve a trade-off, such as potentially offering irrelevant feedback when not needed, which can result in a suboptimal user experience. Therefore, it would be beneficial for users to have control over the level of proactiveness.
Improve trustworthiness of design suggestions
Interestingly, although there is no guarantee that Human experts’ responses are more truthful, participants were more concerned about ChatGPT’s credibility. Future systems take inspiration from the behavior of Human experts in our study. Having a fluid back-and-forth conversation to build a common ground was one thing, but they also often accompanied their responses with rationalized explanations and supplemented supporting empirical evidence. In a similar vein, participants were also uneasy about not being able to verify ChatGPT responses especially when they lacked relevant knowledge. Providing application examples, as they do in practice (Choi et al., 2023b; Kang et al., 2018) might alleviate this concern. These abilities might require more than just prompt engineering, such as augmented retrieval (Lewis et al., 2020).
Integrating into a practitioner’s workflow
Several participants mentioned that they already use ChatGPT for tasks related to writing or debugging code (P10, P12, P13, P16, P18). But, data visualization practice rests on two pillars: design and implementation. Our study sheds light on ChatGPT’s capability in terms of design knowledge. As P13 suggests, the integration of a ChatGPT-like assistant within current tools would be highly beneficial. Such a design assistant could suggest appropriate chart types (P11, P21) or provide rationales or critiques for generated visualizations.
Toward a better design knowledge education
Some participants noted ChatGPT’s potential for education (P20, P21), such as assisting with understanding unfamiliar charts or aiding in writing educational blogs on data visualization. A recent paper identifies challenges in visualization education (Bach et al., 2023), highlighting AI as a valuable tool for formal and informal learning contexts. However, concerns include the consistency and quality of AI-generated content, as differing responses to students could be problematic. There’s also the risk of students over-relying on AI instead of engaging critically with visualizations, e.g., through peer feedback. Future AI systems in education may act as personalized tutors, tracking progress, guiding learning goals, and fostering both comprehension and generation of visualizations.
7. Conclusion
In this study, we explored the potential of ChatGPT as a design companion, focusing on its capacity to provide valuable visualization insights to practitioners. Our findings reveal that ChatGPT capitalizes on a vast knowledge repository to generate diverse and imaginative suggestions. While limitations like the contextual depth of its responses and truthful interaction dynamics exist, participants expressed optimism about its future utility.
For future work, we plan to investigate ways to build a design feedback system based on the lessons learned. This research direction will involve enhancing chart comprehension and enriching the knowledge base of LLMs with well-founded rationales, including visualization examples and empirical evidence, developing user interfaces and interaction designs tailored to the designer’s workflow, and evaluating the system’s practical applicability.
Acknowledgements.
We would like to acknowledge the support of the National Science Foundation (#2146868).
Appendix A Scoring rubrics and guidelines for LLM-based evaluation
In the LLM-based automatic evaluation (Section 5.2), we integrate the scoring rubrics and guidelines listed below as a part of the evaluation prompt. While the scoring rubrics were originally defined in Table 1, we make them more comprehensive and descriptive based on experiments in order to better guide the LLM evaluator.
Table 5. Scoring rubrics for LLM-based evaluation
Metric Score & Description
Coverage 1: Addresses 0–20% of the user’s questions. Almost fails to cover the required points and the majority of the core components, rendering the response ineffective.
2: Addresses 20–40% of the user’s questions. Misses multiple major components of the question, leaving significant portions of the query unanswered.
3: Addresses 40–60% of the user’s questions. Covers the majority of aspects but omits several important components, leaving notable gaps in the response coverage.
4: Addresses 60–80% of the user’s questions. Covers all critical aspects with only minor elements missing, which does not affect the overall completeness of the response.
5: Addresses 80–100% of the user’s questions. Meticulously addresses every aspect of the question with exhaustive detail, ensuring no possible improvement but with no redundant and lengthy description.
Topicality 1: Mostly or entirely off-topic. Fails to maintain any meaningful connection to the subject matter. Major deviation from the question, with significant irrelevant content or a complete failure to address the core question.
2: Partially addresses the topic but frequently digresses. Significant portions stray from the main subject. Noticeable irrelevant points or omissions of several key aspects of the question, leading to a lack of focus.
3: Somewhat relevant; minor deviations from the topic. Contains slight irrelevant content or misses a minor part of the question, but the main focus is generally maintained.
4: Mostly on-topic; occasional tangents. Primarily addresses the question with only a few minor deviations or tangential points, which does not detract from the overall focus.
5: Fully on-topic with no noticeable digressions. Addresses the question with no deviation, and zero irrelevant content. Every sentence directly contributes to answering the question without too much detail.
Breadth 1: Very narrow; offering no alternative ideas. Extremely narrow focus missing most relevant aspects. Shows no awareness of alternative perspectives.
2: Covers a small range of ideas, omitting key concepts. Develops these limited ideas superficially without exploring their connections or implications.
3: Moderately diverse but limited in scope. Addresses some relevant aspects but misses several important areas. Shows restricted perspective with minimal variety.
4: Mostly on-topic; occasional tangents. Primarily addresses the question with only a few minor deviations or tangential points, which does not detract from the overall focus.
5: Fully on-topic with no noticeable digressions. Addresses the question with no deviation, and zero irrelevant content. Every sentence directly contributes to answering the question without too much detail.
Depth 1: Very minimal detail; lacks substance or justification. Offers completely superficial content. Shows no understanding of core concepts or principles.
2: Superficial treatment with limited insight. Provides superficial explanations missing key concepts. Lacks detailed analysis of the concepts presented.
3: Moderately detailed; some analysis or explanation. Presents basic explanations with limited exploration. Provides surface-level analysis without diving into underlying complexities.
4: Good explanation with solid reasoning and insight. Offers solid explanations of core concepts with meaningful connections. Shows clear understanding beyond surface level.
5: Deeper explanation, demonstrated with examples or expertise. Makes sophisticated connections between concepts and demonstrates expert-level response.
Clarity 1: Very unclear; difficult to follow or disorganized. The writing is significantly confusing and often incomprehensible.
2: Somewhat unclear; main ideas partially obscured. The response is poorly organized, with significant logical gaps and a lack of clear structure. Examples, if present, are confusing or irrelevant.
3: Moderately clear; understandable with some effort. The response has noticeable clarity issues, with ideas not flowing smoothly from one to another. Some points require re-reading to grasp their meaning.
4: Clear, well-structured, and easy to read. The response flows smoothly with only occasional minor unclear points. The writing is precise and unambiguous, requiring minimal effort to understand.
5: Flows seamlessly with minimal effort to comprehend. The response is meticulously organized with distinct, logical section breaks. Each section contains precise, directly relevant examples that unequivocally support the points made.
Actionability 1: No actionable guidance or next steps. Completely lacks practical implementation value. Offers no useful guidance or implementable steps. Completely impractical advice.
2: Little actionable guidance; mostly explanations. Provides vague guidance without clear steps. Advice is difficult to implement.
3: Some actionable guidance; partially implementable. Presents some practical advice but with unclear steps. Guidance lacks specificity.
4: Clear, actionable guidance that can be followed easily. Offers practical advice with clear steps, though some details may need clarification.
5: Highly actionable; provides explicit step-by-step instructions. Provides specific, implementable steps with practical examples. Anticipates challenges and offers solutions.
Table 6. Scoring guideline for LLM-based evaluation
Metric Guideline
Coverage - Check if all parts of the question are addressed
- Assess completeness of the response relative to the query
- Look for gaps in addressing key aspects of the question
- Evaluate whether important sub-topics are covered
- Check if the scope of the response matches the question
Topicality - Check if the response directly answers the main question
- Identify any off-topic or tangential information
- Assess whether key aspects of the question are addressed
- Look for focus maintenance throughout the response
- Evaluate relevance of any additional context provided
Breadth - Check coverage of different relevant aspects or perspectives
- Assess variety of examples or use cases provided
- Look for consideration of alternative approaches
- Evaluate inclusion of related concepts or implications
- Check if multiple stakeholder viewpoints are considered
Depth - Assess how thoroughly core concepts are explained
- Look for detailed explanations of underlying principles
- Check for discussion of relationships between concepts
- Evaluate the presence of advanced or nuanced insights
- Consider whether complex aspects are adequately explored
Clarity - Evaluate the logical flow and structure
- Check for clear transitions between ideas
- Assess the use of technical terms and their explanations
- Look for concrete examples that illustrate points
- Check if the response is well-organized with clear paragraphs
Actionability - Assess whether advice can be practically implemented
- Look for specific, concrete steps or recommendations
- Check if prerequisites or requirements are clearly stated
- Evaluate the practicality of suggested solutions
- Consider whether guidance is appropriate for the context